### v1

# Deformations of Anosov subgroups: Limit cones and growth indicators

Abstract.
Let $G$ be a connected semisimple real algebraic group. We prove the continuity of limit cones along deformations of Anosov subgroups of $G$ under a certain convexity assumption. This convexity assumption turns out to be necessary. We discuss an application to the notion of sharpness of an action of a discrete subgroup on a non-Riemannian homogeneous space. We also obtain that growth indicators and certain critical exponents vary continuously in the space of Anosov representations.
1. Introduction
Let $G$ be a connected semisimple real algebraic group. The limit cone of a discrete subgroup of $G$ is a fundamental object in the study of asymptotic properties of discrete subgroups. Let $\mathfrak{a}$ be a Cartan subalgera of the Lie algebra $\mathfrak{a}$ of $G$ and $\mathfrak{a}^{+}\subset\mathfrak{a}$ be a closed positive Weyl chamber. Let $\mu:G\to\mathfrak{a}^{+}$ denote the Cartan projection map (see (2.1)). For any closed subgroup $\Gamma$ of $G$, the limit cone $\mathcal{L}_{\Gamma}$ is the asymptotic cone of the Cartan projection $\mu(\Gamma)$:
$\mathcal{L}_{\Gamma}=\{\lim t_{i}\mu(\gamma_{i})\in\mathfrak{a}^{+}:\gamma_{i}\to\infty\text{ in $\Gamma$},t_{i}>0\}.$
This notion was introduced by Benoist [3]. It is a natural question whether the limit cones of discrete subgroups vary continuously in the deformation space. The main goal of this paper is to answer this question in the affirmative for Anosov subgroups under a certain convexity assumption. This convexity assumption turns out to be necessary. We discuss an application of this result to the notion of sharpness introduced by Kassel and Kobayashi [16]. We also obtain that certain critical exponents and growth indicators vary continuously along deformations of Anosov subgroups.
Anosov subgroups of $G$ are considered as higher rank analogues of convex cocompact subgroups of a rank one Lie group. They play an important role in the study of higher Teichmüller spaces ([21], [12]). Letting $\Pi$ be the set of all simple roots of $G$ with respect to $\mathfrak{a}^{+}$, Anosov subgroups are defined for every non-empty subset $\theta$ of $\Pi$: a finitely generated subgroup $\Gamma<G$ is $\theta$-Anosov if there exists $C>1$ such that for all $\alpha\in\theta$ and $\gamma\in\Gamma$, we have
$\alpha(\mu(\gamma))\geq C^{-1}|\gamma|-C$
where ||$\cdot$|| ⋅ | is the word metric on $\Gamma$ with respect to a fixed finite generating subset. All Anosov subgroups in our paper are assumed to be non-elementary, i.e., virtually non-cyclic.
Limit cones vary continuously
For a finitely generated group $\Gamma$, we denote by $\operatorname{Hom}(\Gamma,G)$ the space of all homomorphisms of $\Gamma$ to $G$. Let $\mathsf{C}(\mathfrak{a}^{+})$ be the space of all closed cones in the positive Weyl chamber $\mathfrak{a}^{+}$. Both have natural topology (see (2.2)). When $\Gamma$ is a subgroup of $G$, we let $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$ be the natural inclusion map from $\Gamma$ to $G$.
Theorem 1.1.
Let $G$ be a connected semisimple real algebraic group. Let $\theta\subset\Pi$ be a non-empty subset. If $\Gamma$ is a $\theta$-Anosov and $\theta$-convex subgroup of $G$ and has reductive [^1] A non-trivial algebraic subgroup of $G$ is reductive if its unipotent radical is trivial Zariski closure, then the map
$\sigma\mapsto\mathcal{L}_{\sigma(\Gamma)}\in\mathsf{C}(\mathfrak{a}^{+})$
is continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
Moreover, the continuity fails in general unless $\Gamma$ is $\theta$-convex.
Let $\mathcal{W}_{\theta}$ be the subgroup consisting of all Weyl elements of $G$ fixing the subspace $\mathfrak{a}_{\theta}=\bigcap_{\alpha\in\Pi-\theta}\operatorname{ker}\alpha$ pointwise. We say that $\Gamma$ is $\theta$-convex if the orbit $\cal W_{\theta\cup\operatorname{i}(\theta)}\mathcal{L}_{\Gamma}$ under the adjoint action is convex in $\mathfrak{a}$, where $\operatorname{i}$ denotes the opposition involution of $G$ (see (2.2)).
The continuity in Theorem 1.1 consists of two parts: lower semicontinuity (Proposition 2.3) and upper semicontinuity (Theorem 4.6). The reductive Zariski closure hypothesis (resp. convexity hypothesis) is not required for the upper (resp. lower) semicontinuity. The failure of continuity in the absence of the convexity assumption is discussed in section 3.
Remark 1.2.
  1. (1)
    If $G$ has rank two, any Anosov subgroup is $\theta$-Anosov and $\theta$-convex for some $\theta\subset\Pi$; hence Theorem 1.1 applies in this case (Corollary 4.10). This is not the case when $\operatorname{rank}G=3$ (see section 3).
  2. (2)
    The continuity of limit cones at $\operatorname{id}_{\Gamma}$ was proved by Kassel [15] when $\Gamma$ is a convex cocompact subgroup of a rank one simple algebraic subgroup of $G$. Indeed, such $\Gamma$ is always a $\theta$-Anosov $\theta$-convex subgroup of $G$ for some $\theta\subset\Pi$ (Lemma 4.8).
  3. (3)
    For a Zariski dense $\Pi$-Anosov subgroup (also known as a Borel-Anosov subgroup), isomorphic to the fundamental group of a closed negatively curved manifold, this was proved by Sambarino using thermodynamic formalism [30].
Our proof of Theorem 1.1 is completely different from [15] and [30], based on the local-to-global principle of Morse quasi-geodesics, due to Kapovich-Leeb-Porti [14].
We discuss some applications of the upper semicontinuity of limit cones (Theorem 4.6).
Sharpness is an open condition
For a closed subgroup $H$ of $G$, a discrete subgroup $\Gamma$ of $G$ is called sharp for $G/H$ if the limit cones of $\Gamma$ and $H$ intersect only at zero:
$\mathcal{L}_{\Gamma}\cap\mathcal{L}_{H}=\{0\}.$
This is an important notion arising in the study of compact Clifford-Klein forms, introduced in [16]. If $\Gamma$ is sharp for $G/H$, then it acts properly discontinuously on $G/H$ ([4], [20]). Kassel and Tholozan proved in their recent work [17] that if $\Gamma$ acts properly discontinuously and cocompactly on $G/H$, then $\Gamma$ is sharp for $G/H$. We have the following immediate consequence of Theorem 4.6 that the sharpness for $G/H$ is an open condition in the following situation:
Corollary 1.3.
Let $H<G$ be a closed subgroup. If $\Gamma$ is a $\theta$-Anosov and $\theta$-convex subgroup of $G$ which is sharp for $G/H$, then there exists an open neighborhood $\cal O$ of $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ such that for all $\sigma\in\cal O$, $\sigma(\Gamma)$ is sharp for $G/H$.
When $H$ is reductive of co-rank one in $G$, this corollary was proved in [17, Corollary 6.3] without the convexity assumption. We refer to Corollary 4.7 for a more general statement which in particular gives another proof of [17, Corollary 6.3].
Growth indicators vary continuously
Our proof of Theorem 1.1 naturally extends to the study of continuity of $\theta$-limit cones (Theorem 5.1). Using this, we show that $\theta$-growth indicators vary continuously in the deformation space of a $\theta$-Anosov subgroup. This can be regarded as a higher rank analog of the classical result that the critical exponents for the Riemannian metric vary continuously (even analytically) in the space of convex cocompact representations of a finitely generated group ([29], [2]).
Let $p_{\theta}:\mathfrak{a}\to\mathfrak{a}_{\theta}$ be the canonical projection invariant under $\mathcal{W}_{\theta}$ where $\mathfrak{a}_{\theta}=\bigcap_{\alpha\in\Pi-\theta}\operatorname{ker}\alpha$. Let $\mathfrak{a}_{\theta}^{+}=\mathfrak{a}_{\theta}\cap\mathfrak{a}^{+}$ and consider the $\theta$-Cartan projection map $\mu_{\theta}=p_{\theta}\circ\mu:G\to\mathfrak{a}_{\theta}^{+}$. For a closed subgroup $\Gamma<G$, its $\theta$-limit cone $\mathcal{L}_{\Gamma}^{\theta}\subset\mathfrak{a}_{\theta}^{+}$ is the asymptotic cone of $\mu_{\theta}(\Gamma)$ and we have
$\mathcal{L}_{\Gamma}^{\theta}=p_{\theta}(\mathcal{L}_{\Gamma}).$
When $\Gamma$ is $\theta$-discrete in the sense that $\mu_{\theta}|_{\Gamma}$ is a proper map, the growth indicator $\psi^{\theta}_{\Gamma}:\mathfrak{a}_{\theta}^{+}\to\mathbb{R}\cup\{-\infty\}$ is defined as follows: if $v\in\mathcal{L}^{\theta}_{\Gamma}$,
$\psi_{\Gamma}^{\theta}(v)=\|v\|\inf_{v\in\cal C}\limsup_{T\to\infty}\frac{1}{T}{\log\#\{\gamma\in\Gamma:\mu_{\theta}(\gamma)\in\cal C,\|\mu_{\theta}(\gamma)\|\leq T\}}$ (1.1)
[…]
For $\theta=\Pi$, any discrete subgroup of $G$ is $\Pi$-discrete and $\psi_{\Gamma}:=\psi_{\Gamma}^{\Pi}$ was introduced by Quint [28]. This notion was extended to a general $\theta$ in [19]. The growth indicator $\psi_{\Gamma}^{\theta}$ describes the exponential growth rate of $\Gamma$ in all directions of $\mathfrak{a}_{\theta}^{+}$, and hence is a higher rank version of the classical Riemannian critical exponent in the rank one group. On the relative interior $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$, $\psi_{\Gamma}^{\theta}$ is strictly positive.
Theorem 1.4.
Let $\Gamma$ be a Zariski dense $\theta$-Anosov subgroup of $G$. If $\sigma_{n}\to\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$, then
$\lim_{n\to\infty}\psi^{\theta}_{\sigma_{n}(\Gamma)}(v)=\psi^{\theta}_{\Gamma}(v)\quad\text{ for all $v\in\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$}$
and the convergence is uniform on compact subsets of $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$.
Using this result, we deduce the continuity of the critical exponents
$\delta_{\Gamma}^{\theta}\coloneqq\limsup_{T\to\infty}\frac{1}{T}{\log\#\{\gamma\in\Gamma:\|\mu_{\theta}(\gamma)\|\leq T\}},$
as well as the maximal growth directions $u_{\Gamma}^{\theta}\in\mathfrak{a}_{\theta}^{+}$ such that $\delta_{\Gamma}^{\theta}=\psi_{\Gamma}^{\theta}(u_{\Gamma}^{\theta})=\max_{v\in\mathfrak{a}_{\theta},\|v\|=1}\psi_{\Gamma}^{\theta}(v)$ ([28], [19]).
Corollary 1.5.
Let $\Gamma$ be a Zariski dense $\theta$-Anosov subgroup of $G$. The maps
$\sigma\mapsto u_{\sigma(\Gamma)}^{\theta}\quad\text{ and }\quad\sigma\mapsto\delta_{\sigma(\Gamma)}^{\theta}$
are continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
For a Borel-Anosov subgroup $\Gamma$, isomorphic to the fundamental group of a closed negatively curved manifold, Theorem 1.4 and Corollary 1.5 were proved by Sambarino [30]. Unlike his proof, our proof uses the conformal measure theory, inspired by the work of Sullivan [32] and McMullen [24].
More general critical exponents
For any discrete subgroup $\Gamma<G$ and a linear form $\phi\in\mathfrak{a}^{*}$ which is positive on $\mathcal{L}_{\Gamma}-\{0\}$, we can define the $\phi$-critical exponent $0\leq\delta_{\phi,\Gamma}<\infty$ by
$\delta_{\phi,\Gamma}=\limsup_{T\to\infty}\frac{\log\#\{\gamma\in\Gamma:\phi(\mu(\gamma))\leq T\}}{T}.$
Theorem 1.6.
Let $\Gamma$ be a $\theta$-Anosov subgroup of $G$ such that $\mathcal{L}_{\Gamma}$ is a convex cone contained in $\mathfrak{a}_{\theta}$. Let $\phi\in\mathfrak{a}^{*}$ be such that $\phi>0$ on $\mathcal{L}_{\Gamma}-\{0\}$. Then for all $\sigma\in\operatorname{Hom}(\Gamma,G)$ sufficiently close to $\operatorname{id}_{\Gamma}$, the critical exponent $0<\delta_{\phi,\sigma(\Gamma)}<\infty$ is well-defined and the map
$\sigma\mapsto\delta_{\phi,\sigma(\Gamma)}$
is continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
Corollary 1.7.
Let $\Gamma$ be a $\theta$-Anosov subgroup of $G$ such that $\mathcal{L}_{\Gamma}$ is a convex cone contained in $\mathfrak{a}_{\theta}$. Let $\phi\in\mathfrak{a}^{*}$ be such that $\phi>0$ on $\mathcal{L}_{\Gamma}-\{0\}$. Suppose that for some sequence $\sigma_{n}\to\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ and $\kappa>0$,
$\sup_{n}\psi_{\sigma_{n}(\Gamma)}(v)\leq\kappa\phi(v)\quad\text{ for all $v\in\mathfrak{a}^{+}$.}$
Then
$\psi_{\Gamma}\leq\kappa\phi.$
Denote by $\rho$ the half sum of all positive roots of $(\mathfrak{g},\mathfrak{a})$ counted with multiplicity. The condition $\psi_{\Gamma}\leq\rho$ is equivalent to the condition that the quasi-regular representation $L^{2}(\Gamma\backslash G)$ is tempered ([10], [23]). Corollary 1.7 applied to the linear form $\rho$ states that if $L^{2}(\sigma_{n}(\Gamma)\backslash G)$ are all tempered, then so is $L^{2}(\Gamma\backslash G)$. This particular fact was proved in [11] via the study of matrix coefficients. But Corollary 1.7 offers a more detailed information on the behavior of growth indicators as we are allowed to use more general linear forms.
Acknowledgement
We would like to thank Fanny Kassel, Misha Kapovich, Dongryul Kim, and Kostas Tsouvalas for useful discussions.
2. Lower semicontinuity of limit cones
Let $G$ be a connected semisimple real algebraic group. Let $A$ be a maximal real split torus of $G$. Let $\mathfrak{g}$ and $\mathfrak{a}$ denote the Lie algebras of $G$ and $A$, respectively. Fix a positive Weyl chamber $\mathfrak{a}^{+}\subset\mathfrak{a}$ and set $A^{+}=\exp\mathfrak{a}^{+}$. Fix a maximal compact subgroup $K<G$ such that the Cartan decomposition $G=KA^{+}K$ holds. For $g\in G$, there exists a unique element $\mu(g)=\mu_{G}(g)\in\mathfrak{a}^{+}$ such that $g\in K\exp(\mu(g))K$, called the Cartan projection of $g$. The Cartan projection map
$\mu=\mu_{G}:G\to\mathfrak{a}^{+}$ (2.1)
is continuous and proper.
[…]
$\|\mu(g)\|=d(go,o)\quad\text{ for all $g\in G$. }$
The Weyl group $\cal W$ is given by $N_{K}(A)/C_{K}(A)$, where $N_{K}(A)$ and $C_{K}(A)$ denote the normalizer and centralizer of $A$ in $K$, respectively. The Weyl group $\cal W$ acts on $\mathfrak{a}$ by adjoint action: $w.v=\operatorname{Ad}_{w}(v)$ for $w\in\cal W$ and $v\in\mathfrak{a}$. Let $\operatorname{i}:\mathfrak{a}\to\mathfrak{a}$ denote the opposition involution, that is,
$\operatorname{i}(v)=-\operatorname{Ad}_{w_{0}}(v)\quad\text{for $v\in\mathfrak{a}$}$ (2.2)
where $w_{0}$ is the longest Weyl element. It induces the involution $\operatorname{i}:\Pi\to\Pi$ which we also denote by $\operatorname{i}$: $\operatorname{i}(\alpha)=\alpha\circ\operatorname{i}$ for all $\alpha\in\Pi$.
Limit cone
The limit cone $\mathcal{L}_{\Gamma}=\mathcal{L}_{\Gamma,G}$ of a closed subgroup $\Gamma$ of $G$ is defined as the asymptotic cone of $\mu(\Gamma)$:
$\mathcal{L}_{\Gamma}=\{\lim t_{i}\mu(\gamma_{i})\in\mathfrak{a}^{+}:\gamma_{i}\to\infty\text{ in $\Gamma$},t_{i}>0\}.$
Since $\mu(g^{-1})=\operatorname{i}(\mu(g))$ for all $g\in G$, it follows that $\mathcal{L}_{\Gamma}$ is invariant under the opposition involution $\operatorname{i}$.
Any $g\in G$ can be written as the commuting product $g=g_{h}g_{e}g_{u}$ where $g_{h}$ is hyperbolic, $g_{e}$ is elliptic and $g_{u}$ is unipotent. The hyperbolic component $g_{h}$ is conjugate to a unique element of the form $\exp(\lambda(g))$ where $\lambda(g)\in\mathfrak{a}^{+}$. The element $\lambda(g)$ is called the Jordan projection of $g$. The Jordan projection $\lambda:G\to\mathfrak{a}^{+}$ is continuous.
For a closed subgroup $\Gamma$ of $G$, we denote by $\mathcal{L}_{\Gamma}^{\operatorname{Jor}}=\mathcal{L}_{\Gamma,G}^{\operatorname{Jor}}$ the smallest closed cone of $\mathfrak{a}^{+}$ containing the Jordan projection $\lambda(\Gamma)$. The relationship between $\mathcal{L}_{\Gamma}$ and $\mathcal{L}_{\Gamma}^{\operatorname{Jor}}$ will be important to our proofs. Since for any $g\in G$,
$\lambda(g)=\lim_{n\to\infty}\frac{\mu(g^{n})}{n},$
we have the inclusion
$\mathcal{L}_{\Gamma}\supset\mathcal{L}_{\Gamma}^{\rm Jor}$ (2.3)
for any discrete subgroup $\Gamma$ of $G$. In general, $\cal L_{\Gamma}\neq\mathcal{L}_{\Gamma}^{\rm Jor}$. For example, when $\Gamma$ is a unipotent subgroup, $\mathcal{L}_{\Gamma}^{\rm Jor}=\{0\}$ while $\cal L_{\Gamma}$ is non-trivial. However, if $\Gamma$ has reductive Zariski closure in $G$, these two are equal:
Theorem 2.1.
Let $\Gamma$ be a closed subgroup whose Zariski closure is reductive. Then
$\mathcal{L}_{\Gamma}=\mathcal{L}_{\Gamma}^{\operatorname{Jor}}.$ (2.4)
Moreover, if $\Gamma$ is Zariski dense in $G$, $\mathcal{L}_{\Gamma}$ is convex and has non-empty interior.
Proof.
This was proved by Benoist in [3] when $G$ is reductive and $\Gamma$ is Zariski dense in $G$. The general case can be deduced as follows. If the Zariski closure of $\Gamma$, say, $H$, is reductive, then we have a compatible Cartan decomposition $H=(K\cap H)(A\cap H)(K\cap H)$ by replacing $K$ and $A$ up to conjugation if necessary [25]. If $\mathfrak{a}_{H}^{+}$ is a positive Weyl chamber of $\operatorname{Lie}(A\cap H)$, we have
$\mathcal{L}_{\Gamma,H}=\mathcal{L}^{\operatorname{Jor}}_{\Gamma,H}$
by [3]. Since $\mathfrak{a}_{H}^{+}\subset\mathfrak{a}$, any $v\in\mathfrak{a}_{H}^{+}$ lies in $w^{-1}.\mathfrak{a}^{+}$ for some $w\in\cal W$ and hence $w.v=\mu_{G}(\exp v)$. Suppose that $v=\lim t_{i}\mu_{H}(\gamma_{i})\in\mathcal{L}_{\Gamma,H}\cap w^{-1}.\operatorname{int}\mathfrak{a}^{+}$ for some $w\in\mathcal{W}$. Then $\mu_{H}(\gamma_{i})\in w^{-1}.\operatorname{int}\mathfrak{a}^{+}$ for all sufficiently large $i$ and hence $w.\mu_{H}(\gamma_{i})=\mu_{G}(\gamma_{i})$. So
$\mu_{G}(\exp v)=w.v=\lim t_{i}w.\mu_{H}(\gamma_{i})=\lim t_{i}\mu_{G}(\gamma_{i}).$
By the continuity of the Cartan projection $\mu$, it follows that
$\mathcal{L}_{\Gamma,G}=\mu_{G}(\exp\mathcal{L}_{\Gamma,H}).$
Similarly, we have $\mathcal{L}^{\operatorname{Jor}}_{\Gamma,G}=\mu_{G}(\exp\mathcal{L}_{\Gamma,H}^{\operatorname{Jor}})$. Therefore (2.4) follows. ∎
Definition 2.2.
  • For a finitely generated group $\Gamma$, let $\operatorname{Hom}(\Gamma,G)$ be the space of all homomorphisms of $\Gamma$ to $G$. The topology on $\operatorname{Hom}(\Gamma,G)$ is given as follows: for a fixed finite generating subset $S$ of $\Gamma$, a sequence $\sigma_{n}\in\operatorname{Hom}(\Gamma,G)$ converges to $\sigma\in\operatorname{Hom}(\Gamma,G)$ as $n\to\infty$ if for each $\gamma\in S$, $\sigma_{n}(\gamma)\to\sigma(\gamma)$ as $n\to\infty$. This topology is independent of the choice of a generating subset.
  • We denote by $\mathsf{C}(\mathfrak{a}^{+})$ the space of all (non-zero) closed cones of $\mathfrak{a}^{+}$ equipped with the topology: $\mathcal{C}_{i}\to\mathcal{C}$ if the Hausdorff distance between $\mathbb{P}(\mathcal{C}_{i})$ and $\mathbb{P}(\mathcal{C})$ in $\mathbb{P}(\mathfrak{a}^{+})$ converges to zero as $i\to\infty$. Under this topology, ${\mathsf{C}}(\mathfrak{a}^{+})$ is a compact metrizable space.
Throughout the paper, all closed cones in $\mathfrak{a}^{+}$ are assumed to be non-degenerate, i.e., they are not $\{0\}$.
We prove the lower semicontinutity of limit cones under the reductive Zariski closure hypothesis:
Proposition 2.3 (Lower semicontinuity).
Let $\Gamma$ be a finitely generated discrete subgroup of $G$. If $\sigma_{n}$ is a sequence of discrete representations of $\Gamma$ converging to $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$, then any accumulation point $\mathcal{L}$ of the sequence $\mathcal{L}_{\sigma_{n}(\Gamma)}$ in $\mathsf{C}(\mathfrak{a}^{+})$ contains $\mathcal{L}^{\operatorname{Jor}}_{\Gamma}$.
In particular, if $\mathcal{L}_{\Gamma}=\mathcal{L}_{\Gamma}^{\operatorname{Jor}}$ (e.g., the Zariski closure of $\Gamma$ is reductive), then $\mathcal{L}$ contains $\mathcal{L}_{\Gamma}$.
Proof.
By passing to a subsequence, we may assume that $\mathcal{L}_{\sigma_{n}(\Gamma)}\to\mathcal{L}$ as $n\to\infty$. Since $\mathbb{P}(\mathcal{L}_{\Gamma})$ is compact, it is enough to show that for all unit vectors $v\in\mathcal{L}_{\Gamma}$, the distance $d_{\mathfrak{a}}(v,\mathcal{L}_{\sigma_{n}(\Gamma)})$ converges to 00 as $n\to\infty$.
Let $v\in\mathcal{L}^{\operatorname{Jor}}_{\Gamma}$ be a unit vector. Then there exists a sequence $\gamma_{k}\in\Gamma$ such that
$\|\lambda^{\prime}(\gamma_{k})-v\|<\frac{1}{k}\quad\text{for all $k\geq 1$,}$
where $\lambda^{\prime}(\gamma_{k})=\|\lambda(\gamma_{k})\|^{-1}\lambda(\gamma_{k})$ denotes the unit vector in the direction of $\lambda(\gamma_{k})$. Since Jordan projection varies continuously, for each $k$, we can choose $n_{k}$ large enough such that $\|\lambda^{\prime}(\gamma_{k})-\lambda^{\prime}(\sigma_{n_{k}}(\gamma_{k}))\|<\frac{1}{k}$. Thus
$\|\lambda^{\prime}(\sigma_{n_{k}}(\gamma_{k}))-v\|<\frac{2}{k}\quad\text{ for all $k$.}$
Since $\lambda^{\prime}(\sigma_{n_{k}}(\gamma_{k}))\in\mathcal{L}^{\operatorname{Jor}}_{\sigma_{n_{k}}(\Gamma)}$, this proves that
$d_{\mathfrak{a}}(\mathcal{L}^{\operatorname{Jor}}_{\sigma_{n_{k}}(\Gamma)},v)\to 0\quad\text{as }n\to\infty.$ (2.5)
On the other hand, by (2.3), we have $\mathcal{L}_{\sigma_{n_{k}}(\Gamma)}\supset\mathcal{L}^{\operatorname{Jor}}_{\sigma_{n_{k}}(\Gamma)}$. Thus (2.5) implies that
$d_{\mathfrak{a}}(\mathcal{L}_{\sigma_{n_{k}}(\Gamma)},v)\to 0\quad\text{as }n\to\infty,$
showing that $v\in\mathcal{L}$. Therefore $\mathcal{L}\supset\mathcal{L}^{\operatorname{Jor}}_{\Gamma}$. Hence this finishes the proof together with Theorem 2.1. ∎
Since $\mu(\Gamma)$ is not contained in $\mathcal{L}_{\Gamma}$ in general, one cannot replace the Jordan projection $\lambda(\gamma_{k})$ by the Cartan projection $\mu(\gamma_{k})$ in the above proof.
3. Failure of the continuity of limit cones: an example
Upper semicontinuity of limit cones does not hold in general, even for the class of Anosov subgroups. In this section, we present an example illustrating this point. These examples are $\theta$-Anosov subgroups of $\operatorname{SL}(4,\mathbb{R})$ but are not $\theta$-convex.
We begin by defining $\theta$-Anosov subgroups and the $\theta$-convexity condition. Let $G$ be a connected semisimple real algebraic group. We keep the notation set up in section 2. Hence $\Pi$ denotes the set of simple roots of $(\mathfrak{g},\mathfrak{a})$.
Definition 3.1.
For a finitely generated subgroup $\Gamma<G$ and a non-empty subset $\theta\subset\Pi$, a representation $\sigma:\Gamma\to G$ is called $\theta$-Anosov if there exists a constant $C>1$ such that for all $\alpha\in\theta$ and $\gamma\in\Gamma$,
$\alpha(\mu(\sigma(\gamma))\geq C^{-1}|\gamma|-C$ (3.1)
where ||$\cdot$|| ⋅ | denotes the word metric on $\Gamma$ with respect to a fixed finite generating subset of $\Gamma$. We say that $\Gamma$ is a $\theta$-Anosov subgroup of $G$ if the natural inclusion map $\operatorname{id}_{\Gamma}$ is $\theta$-Anosov.
A $\Pi$-Anosov subgroup is also called a Borel-Anosov subgroup. By an Anosov subgroup, we mean a $\theta$-Anosov subgroup for some non-empty $\theta\subset\Pi$. As mentioned in the introduction, all Anosov subgroups in our paper are assumed to be non-elementary, i.e., not virtually cyclic.
Let $\theta$ be a non-empty subset of $\Pi$. Each $\theta$-Anosov representation is discrete and has finite kernel [12]. Since $\mu(g)=\operatorname{i}\mu(g^{-1})$ for all $g\in G$, the $\theta$-Anosov condition is same as the $\theta\cup\operatorname{i}(\theta)$-Anosov condition. It is immediate that for a $\theta$-Anosov subgroup $\Gamma<G$, its limit cone $\mathcal{L}_{\Gamma}$ is disjoint from $\ker\alpha-\{0\}$ for all $\alpha\in\theta$.
Definition 3.2.
Set
$\mathfrak{a}_{\theta}=\bigcap_{\alpha\in\Pi-\theta}\ker\alpha.$ (3.2)
Denote by $\cal W_{\theta}$ the subgroup consisting of all elements of the Weyl group $\cal W$ fixing $\mathfrak{a}_{\theta}$ pointwise; it is generated by reflections with respect to the walls $\ker\alpha$, $\alpha\in\Pi-\theta$. A discrete subgroup $\Gamma$ of $G$ is called $\theta$-convex if the orbit
$\mathcal{W}_{\theta\cup\operatorname{i}(\theta)}\mathcal{L}_{\Gamma}=\bigcup_{w\in\cal W_{\theta\cup\operatorname{i}(\theta)}}\operatorname{Ad}_{w}\mathcal{L}_{\Gamma}\text{ is a convex subset of }\mathfrak{a}.$
Since $\mathcal{W}_{\theta\cup\operatorname{i}(\theta)}\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{+}=\mathcal{L}_{\Gamma}$, the $\theta$-convexity of $\Gamma$ implies that the limit cone $\mathcal{L}_{\Gamma}$ is convex. In the rest of this section, we set
$G=\operatorname{SL}(4,\mathbb{R}).$
We fix the positive Weyl chamber $\mathfrak{a}^{+}$ of $G$ as follows:
$\mathfrak{a}^{+}=\{v=\text{diag}(v_{1},v_{2},v_{3},v_{4}):v_{1}\geq v_{2}\geq v_{3}\geq v_{4},\;\sum_{i=1}^{4}v_{i}=0\}.$
For simplicity, we identify $v=\text{diag}(v_{1},v_{2},v_{3},v_{4})$ with $v=(v_{1},v_{2},v_{3},v_{4})$. The set $\Pi$ of simple roots is given by
$\alpha_{i}(v)=v_{i}-v_{i+1}\quad\text{for $1\leq i\leq 3$}.$
Consider the following block diagonal subgroup:
$H=\begin{pmatrix}\operatorname{SL}(3,\mathbb{R})&0\\ 0&1\end{pmatrix}<\operatorname{SL}(4,\mathbb{R}).$
Observe that the Cartan projection $\mu(H)=\mu_{G}(H)\subset\mathfrak{a}^{+}$ is a folded plane (see Figure 1):
$\mu(H)=V_{1}\cup V_{2}$
where
$V_{1}=\{(v_{1},v_{2},0,v_{3})\in\mathfrak{a}^{+}:v_{1}\geq v_{2}\geq 0\}$
and
$V_{2}=\{(v_{1},0,v_{2},v_{3})\in\mathfrak{a}^{+}:v_{1}\geq 0\geq v_{2}\}.$
$\operatorname{ker}\alpha_{2}$$\operatorname{ker}\alpha_{1}$$\operatorname{ker}\alpha_{3}$
Figure 1. The folded “plane” in the unit sphere of $\mathfrak{a}^{+}$
(green), containing the limit cone (darker green)
More precisely, consider the Cartan projection of $H$: $\mu_{H}:H\to\mathfrak{a}_{H}^{+}$ where
$\mathfrak{a}_{H}^{+}=\{(v_{1},v_{2},v_{3},0):v_{1}\geq v_{2}\geq v_{3},v_{1}+v_{2}+v_{3}=0\}.$
If $\mu_{H}(h)=(v_{1},v_{2},v_{3},0)$ for $h\in H$, then $\mu(h)\in V_{1}\cup V_{2}$ is given by
$\mu(h)=\begin{cases}(v_{1},v_{2},0,v_{3})&\text{ if $v_{2}\geq 0$}\\ (v_{1},0,v_{2},v_{3})&\text{ if $v_{2}\leq 0$}.\end{cases}$
Also consider the ray
$V_{0}=V_{1}\cap V_{2}=\{(v_{1},0,0,-v_{1}):v_{1}\geq 0\}.$
Lemma 3.3.
Let $\theta=\{\alpha_{1},\alpha_{3}\}$. Any Zariski dense Borel-Anosov subgroup $\Gamma$ of $H$ is a $\theta$-Anosov subgroup of $G$ which is not $\theta$-convex. Moreover, for each $i=1,2$,
$\mathcal{L}_{\Gamma}\cap V_{i}\not\subset V_{0}.$ (3.3)
Proof.
Fix a word metric ||$\cdot$|| ⋅ | on $\Gamma$ with respect to a finite generating subset of $\Gamma$. Since $\Gamma$ is a Borel-Anosov subgroup of $H$, there exists $C>1$ such that for any $\gamma\in\Gamma$,
$\alpha_{1}(\mu_{H}(\gamma))\geq C^{-1}|\gamma|-C.$
For $\gamma\in\Gamma$ with $\mu_{H}(\gamma)=(v_{1},v_{2},v_{3},0)$, we then have $\alpha_{1}(\mu(\gamma))=v_{1}-v_{2}$ if $\mu(\gamma)\in V_{1}$ or $\alpha_{1}(\mu(\gamma))=v_{1}$ if $\mu(\gamma)\in V_{2}$. In the second case, we have $0\geq v_{2}\geq v_{3}=-v_{1}-v_{2}$, and hence $v_{1}\geq-2v_{2}\geq-v_{2}$. So $v_{1}\geq\frac{1}{2}(v_{1}-v_{2})$. Therefore for any $\gamma\in\Gamma$,
$\alpha_{1}(\mu(\gamma))\geq\frac{1}{2}\alpha_{1}(\mu_{H}(\gamma))\geq\frac{1}{2C}|\gamma|-2C.$
It follows that $\Gamma$ is an $\{\alpha_{1}\}$-Anosov subgroup of $G$ or, equivalently, an $\{\alpha_{1},\alpha_{3}\}$-Anosov subgroup of $G$ since $\operatorname{i}(\alpha_{1})=\alpha_{3}$.
We now claim that $\Gamma$ is not $\theta$-convex for $\theta=\{\alpha_{1},\alpha_{3}\}$. Since $\Gamma$ is a Zariski dense Borel-Anosov subgroup of $H$, $\mathcal{L}_{\Gamma,H}$ is a convex cone with non-empty interior and contained in $\operatorname{int}\mathfrak{a}_{H}^{+}\cup\{0\}$. Hence $\mathcal{L}_{\Gamma}-\{0\}$ is contained in the union of relative interiors of $V_{1}$ and $V_{2}$, $\mathcal{L}_{\Gamma}\cap V_{i}\not\subset V_{0}$ and $\mathcal{L}_{\Gamma}$ intersects each relative interior of $V_{i}$ non-trivially. Hence $\mathcal{L}_{\Gamma}$ is not convex, and as a consequence, $\Gamma$ is not $\theta$-convex. ∎
Let $\Gamma$ be a Zariski-dense Schottky subgroup of $H$ generated by two diagonalizable elements
$a=\exp v=\text{diag}(e^{v_{1}},e^{v_{2}},e^{v_{3}},1)$
and
$b=g\text{diag}(e^{w_{1}},1,e^{-w_{1}},1)g^{-1}\quad\text{for some $g\in H$}$
where $v_{1}>v_{2}>0>v_{3}$ and $w_{1}>0$. We refer to [9] for a precise definition of a higher rank Schottky subgroup and for the proof that a Schottky subgroup is a Borel-Anosov subgroup.
Proposition 3.4.
For the above Zariski dense Schottky subgroup $\Gamma<H$, the map
$\sigma\mapsto\mathcal{L}_{\sigma(\Gamma)}$
is not continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
Proof.
Since the diagonal entries of $a=\exp v$ are all distinct, $a$ is a loxodromic element of $G=\operatorname{SL}(4,\mathbb{R})$. It follows from [33, Proposition 4.4] that the union of all Zariski-closed and Zariski-connected proper subgroups of $G$ containing $a$ is contained in a proper Zariski-closed subset of $G$, say $\cal Z$. Since $\Gamma$ is not Zariski dense in $G$, we have
$b\in\cal Z.$
Since $G-\cal Z$ is Zariski open and the set of all loxodromic elements of $G$ is Zariski dense ([26], [3, Corollary 1]), we can find a sequence $b_{n}\in G-\cal Z$ of loxodromic elements converging to $b$.
Since $\Gamma$ is a free group generated by $a$ and $b$, we can consider the representation $\sigma_{n}:\Gamma\to G$ given by
$\sigma_{n}(a)=a\quad\text{and}\quad\sigma_{n}(b)=b_{n}.$
Then $\sigma_{n}\to\operatorname{i}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ as $n\to\infty$. Let $\theta=\{\alpha_{1},\alpha_{3}\}$. Since $\Gamma$ is $\theta$-Anosov by Lemma 3.3 and the set of $\theta$-Anosov representations is open in $\operatorname{Hom}(\Gamma,G)$ [12], each $\sigma_{n}(\Gamma)$ is a $\theta$-Anosov subgroup, and hence a discrete subgroup of $G$ for all sufficiently large $n$.
Since $\sigma_{n}(\Gamma)$ is Zariski dense in $G$ (as $b_{n}\notin\cal Z$), the limit cone $\mathcal{L}_{\sigma_{n}(\Gamma)}$ must be a convex cone of $\mathfrak{a}^{+}$. We claim that the upper-semicontinuity of the limit cones fail: we can find an open cone $\cal C\subset\mathfrak{a}^{+}$ containing $\mathcal{L}_{\Gamma}-\{0\}$ which does not contain $\mathcal{L}_{\sigma_{n}(\Gamma)}-\{0\}$ for any large enough $n$.
Consider two open cones $\cal C_{1}$ and $\cal C_{2}$ in $\mathfrak{a}^{+}$ containing $V_{1}-\{0\}$ and $V_{2}-\{0\}$ respectively and $\cal C_{1}\cap\cal C_{2}=V_{1}\cap V_{2}-\{0\}=V_{0}-\{0\}$. By taking $\cal C_{1}$ and $\cal C_{2}$ small enough, we can assume that for any unit vectors $w_{1}\in\cal C_{1}-V_{0}$ and $w_{2}\in\cal C_{2}-V_{0}$, the convex combination $\frac{1}{2}(w_{1}+w_{2})$ does not belong to $\cal C_{1}\cup\cal C_{2}$. Set $\cal C=\cal C_{1}\cup\cal C_{2}$. Since
$\mathcal{L}_{\Gamma}\cap V_{i}\not\subset V_{0}$
by Lemma 3.3, there must be a sequence of unit vectors $u_{n}\in\mathcal{L}_{\sigma_{n}(\Gamma)}$ converging to some unit vector $u\in V_{2}-V_{0}$. By the choice of $a$, we have $\mu(a)=(v_{1},v_{2},0,v_{3})\in V_{1}-V_{0}$. Since $\mu(a)\in V_{1}-V_{0}\subset\cal C_{1}-V_{0}$ and $u_{n}\in V_{2}-V_{0}\subset\cal C_{2}-V_{0}$ for all sufficiently large $n$, we have
$w_{n}\coloneqq\frac{1}{2}(u_{n}+\|\mu(a)\|^{-1}\mu(a))\not\in\cal C$
for all large $n$. On the other hand, $w_{n}$ must belong to in $\mathcal{L}_{\sigma_{n}(\Gamma)}$ by the convexity of $\mathcal{L}_{\sigma_{n}(\Gamma)}$. This shows that $\mathcal{L}_{\sigma_{n}(\Gamma)}-\{0\}\not\subset\cal C$ for all large enough $n$. This finishes the proof. ∎
Note that $\Gamma$ is not Zariski dense in $\operatorname{SL}(4,\mathbb{R})$ in the above example. We remark that Danciger-Guéritaud-Kassel announced that there exists a Zariski dense Anosov subgroup of $\operatorname{PSL}(2,\mathbb{R})\times\operatorname{PSL}(2,\mathbb{R})\times\operatorname{PSL}(2,\mathbb{R})$ where the limit cone does not vary continuously.
4. Upper semicontinuity of limit cones
Let $G$ be a connected semisimple real algebraic group. In this section, we prove the upper semicontinuity of the limit cones of $\theta$-Anosov and $\theta$-convex subgroups of $G$. By the definition of $\theta$-Anosov and $\theta$-convex condition, we may assume without loss of generality that $\theta$ is $\operatorname{i}$-invariant, which we are fixing throughout this section.
The main tool of our proof is the local-to-global principle for Morse quasigeodesics due to Kapovich-Leeb-Porti ([14, Theorem 1.1], see Theorem 4.3). In order to explain this principle, we need to recall some terminology. Recall the notation $\mathfrak{a}_{\theta}=\bigcap_{\alpha\in\Pi-\theta}\ker\alpha$ and $\cal W_{\theta}$ from Definition 3.2. Let $M_{\theta}$ denote the centralizer of $\exp\mathfrak{a}_{\theta}$ in $K$.
The notion of a $\theta$-admissible cone is crucial:
Definition 4.1.
A closed cone $\cal C\subset\mathfrak{a}^{+}$ is called $\theta$-admissible if
  1. (1)
    $\cal C$ is $\operatorname{i}$-invariant: $\operatorname{i}(\cal C)=\cal C$;
  2. (2)
    $\cal C$ is $\theta$-convex, i.e., $\cal W_{\theta}.\cal C$ is convex;
  3. (3)
    $\cal C\cap\left(\bigcup_{\alpha\in\theta}\ker\alpha\right)=\{0\}$.
Note that if $\Gamma$ is $\theta$-Anosov and $\theta$-convex, then its limit cone $\mathcal{L}_{\Gamma}$ is $\theta$-admissible.
Local-to-global principle
Let $\cal C$ be a $\theta$-admissible cone of $\mathfrak{a}^{+}$. Recall $X=G/K$ and $o=[K]\in X$. Let $x$ and $y$ be a $\cal C$-regular pair in $X$, that is,
$\mu(g_{1}^{-1}g_{2})\in\cal C$
for any $g_{1},g_{2}\in G$ such that $x=g_{1}o$ and $y=g_{2}o$. The $\cal C$-cone emanating from $x$ and passing through $y$ is defined as
[…]
The $\cal C$-diamond connecting $x$ to $y$ is defined as
$\Diamond C(x,y)=VC(x,y)\cap VC(y,x).$ (4.1)
Cones and diamonds can be viewed as higher rank generalizations of geodesic rays and segments in the hyperbolic space.
Definition 4.2.
Let $I\subset\mathbb{R}$ be an interval, and $L\geq 1$ and $D\geq 0$ be constants.
  1. (1)
    A map $f:I\to X$ is called $(\cal C,D,L)$-Morse, if $f$ is an $L$-quasi-isometric embedding [^2] That is, $L^{-1}|t-s|-L\leq d(f(t),f(s))\leq L|t-s|+L$ for all $t,s\in I$ and for all $s\leq t$ in $I$, the image $f([s,t])$ lies in the $D$-neighborhood of some $\cal C$-diamond $\Diamond_{\cal C}(x,y)$ in $X$, where the tips $x$ and $y$ are a $\cal C$-regular pair such that $d(x,f(s))\leq D$ and $d(y,f(t))\leq D$.
  2. (2)
    For $S\geq 1$, a map $f:I\to X$ is called $(\cal C,D,L,S)$-local Morse if for all $s,t\in I$ with $0\leq t-s\leq S$, the restriction map $f|_{[s,t]}$ is $(\cal C,D,L)$-Morse.
  3. (3)
    More generally, for a geodesic Gromov hyperbolic space $Y$, a map $f:Y\to X$ is called $(\cal C,D,L)$-Morse (resp. $(\cal C,D,L,S)$-local Morse) if the restriction of $f$ to each unit-speed parametrized geodesic in $Y$ is $(\cal C,D,L)$-Morse (resp. $(\cal C,D,L,S)$-local Morse).
We remark that the Morse datum is invariant under post composition by any element $g\in G$, considered as an isometry of $X$. More precisely, if $f:Y\to X$ is $(\cal C,D,L)$-Morse (resp. $(\cal C,D,L,S)$-local Morse), then so is $g\circ f$ for any $g\in G$.
Finally, we can state the local-to-global principle for Morse quasigeodesics:
Theorem 4.3 (Kapovich-Leeb-Porti, [14, Theorem 1.1]).
Let $\cal C,\cal C^{\prime}\subset\mathfrak{a}^{+}$ be $\theta$-admissible closed cones with nonempty interior such that $\cal C-\{0\}\subset{\rm int}\,\cal C^{\prime}$. For any interval $I\subset\mathbb{R}$ and any constants $L\geq 1$ and $D\geq 0$, there exist $L^{\prime},S\geq 1$ and $D^{\prime}\geq 0$ such that every $(\cal C,D,L,S)$-local Morse map
$f:I\to X$
[…]
Upper semicontinuity of limit cone
In order to apply Theorem 4.3 in our setting, we will need the following lemma on the existence of $\theta$-admissible cones:
Lemma 4.4.
Let $\mathcal{D}$ be a $\theta$-admissible cone of $\mathfrak{a}^{+}$. Then for any open cone $\cal C_{0}$ in $\mathfrak{a}^{+}$ containing $\mathcal{D}-\{0\}$, there exists a $\theta$-admissible closed cone $\cal C$ such that
$\cal D-\{0\}\subset\operatorname{int}\cal C\subset\cal C\subset\cal C_{0}\cup\{0\}.$
Proof.
We observe that $\cal W_{\theta}{}\mathfrak{a}^{+}$ is equal to the union of all Weyl chambers containing $\mathfrak{a}_{\theta}^{+}$, and hence is a convex cone.
Let $\alpha\in\theta$. Since $\ker\alpha\cap\mathfrak{a}^{+}$ is contained in the boundary of $\cal W_{\theta}{}\mathfrak{a}^{+}$, it follows from the convexity of $\cal W_{\theta}{}\mathfrak{a}^{+}$ that $\cal W_{\theta}{}\mathfrak{a}^{+}$ is contained in the half space $\{\alpha\geq 0\}.$ Hence both $\operatorname{int}\mathfrak{a}_{\theta}^{+}$ and $\cal W_{\theta}{}\cal D-\{0\}$ are contained in the open half-space $\{\alpha>0\}$. Therefore we can find a linear form $h_{\alpha}\in\mathfrak{a}^{*}$ such that
$(\ker\alpha\cap\mathfrak{a}^{+})-\{0\}\subset\{h_{\alpha}<0\}\quad\text{and}\quad(\cal C_{0}\cup\cal W_{\theta}{}\cal D)-\{0\}\subset\{h_{\alpha}>0\}.$
Now set
$H\coloneqq\bigcap_{\alpha\in\theta,w\in\cal W_{\theta}}\{h_{\alpha}\circ\operatorname{Ad}_{w}\geq 0\},$ (4.2)
which is clearly a $\cal W_{\theta}$-invariant convex cone. By our choice of $h_{\alpha}$, the interior $\operatorname{int}H$ contains $\cal D-\{0\}$. Since $\theta=\operatorname{i}(\theta)$, we have that $\operatorname{i}(H)$ is also a $\cal W_{\theta}$-invariant convex cone whose interior contains $\cal D-\{0\}=\operatorname{i}(\cal D)-\{0\}$.
Since $\cal D-\{0\}\subset\cal C_{0}$, we can find $\varepsilon>0$ such that the cone $\cal D_{\varepsilon}:=\mathbb{R}_{+}(\cal N_{\varepsilon}(\mathfrak{a}^{1}\cap\cal D))$ is contained in $\cal C_{0}\cup\{0\}$ where $\mathfrak{a}^{1}=\{\|w\|=1\}$ is the unit sphere and $\cal N_{\varepsilon}(\mathfrak{a}^{1}\cap\cal D)=\{w\in\mathfrak{a}^{1}:\|w-(\mathfrak{a}^{1}\cap\cal D)\|\leq\varepsilon\}$.
Define
[…]
By construction, we have $\cal C\cap\left(\bigcup_{\alpha\in\theta}\ker\alpha\right)=\{0\}$. Moreover, $\cal C$ is a closed cone in $\mathfrak{a}^{+}$ whose interior contains $\cal D-\{0\}$. Since $\operatorname{i}(\cal D)=\cal D$ and $\operatorname{i}$ preserves the norm on $\mathfrak{a}$, we have $\operatorname{i}(\cal C)=\cal C$. Since $\mathcal{W}_{\theta}$ acts by isometries on $\mathfrak{a}^{1}$, we have $\mathcal{W}_{\theta}(\cal D_{\varepsilon})=\mathbb{R}_{+}\cal N_{\varepsilon}(\cal W_{\theta}\mathfrak{a}^{1}\cap\cal W_{\theta}\cal D)$. By the hypothesis that $\mathcal{W}_{\theta}\cal D$ is convex, $\mathcal{W}_{\theta}(\cal D_{\varepsilon})$ is convex as well. Since
[…]
it follows that $\cal W_{\theta}{}\cal C$ is convex. Therefore $\cal C$ is $\theta$-admissible. ∎
Proposition 4.5.
Let $\Gamma$ be a $\theta$-Anosov subgroup of $G$. Let $\cal D\subset\mathfrak{a}^{+}$ be a $\theta$-admissible cone containing $\mathcal{L}_{\Gamma}$. Then for any open cone $\cal C_{0}\subset\mathfrak{a}^{+}$ containing $\cal D-\{0\}$, there exists an open neighborhood $\cal O\subset\operatorname{Hom}(\Gamma,G)$ of $\operatorname{id}_{\Gamma}$ such that for all $\sigma\in\cal O$, we have
$\mathcal{L}_{\sigma(\Gamma)}-\{0\}\subset\cal C_{0}.$
Proof.
By Lemma 4.4, we can choose $\theta$-admissible closed cones $\cal C$ and $\cal C^{\prime}$ contained in $\cal C_{0}\cup\{0\}$ so that
$\cal D-\{0\}\subset\operatorname{int}\cal C\subset\cal C\subset\operatorname{int}\cal C^{\prime}\cup\{0\}.$
Consider a left-invariant word metric ||$\cdot$|| ⋅ | on $\Gamma$ with respect to a fixed symmetric finite generating set for $\Gamma$. Since $\Gamma$ is $\theta$-Anosov, there exist constants $L,D\geq 1$ such that the orbit map
$\Gamma\to X$, $\gamma\mapsto\gamma o$
[…]
Denote by $\mathfrak{o}_{n}:\Gamma\to X$ the orbit map of $\Gamma$ via $\sigma_{n}$:
$\gamma\mapsto\sigma_{n}(\gamma)o.$
[…]
Proof of claim.
Consider the $S$-ball $B_{\Gamma}(e,S)\subset\Gamma$ with respect to the word metric centered at the identity element $e$. Let $f:I\cap\mathbb{Z}\to B_{\Gamma}(e,S)$ be a geodesic, where $I\subset\mathbb{R}$ is an interval. Since $\sigma_{n}$ converge to $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$, there exists $n_{0}\geq 1$ such that for all $n\geq n_{0}$,
$\max_{\gamma\in B_{\Gamma}(e,S)}d(\gamma o,\sigma_{n}(\gamma)o)\leq\frac{1}{2}.$
Since the orbit map $\gamma\mapsto\gamma o$ is $(\cal C,D,L)$-Morse, this implies that the restriction $\mathfrak{o}_{n}$ to $B_{\Gamma}(e,S)$ is $(\cal C,D+1,L+1)$-Morse for all $n\geq n_{0}$.
It follows that for all $n\geq n_{0}$, the map $\mathfrak{o}_{n}$ is $(\cal C,D+1,L+1,S)$-local Morse. Indeed, let $s\leq t$ be such that $t-s\leq S$, and let $f:[s,t]\cap\mathbb{Z}\to\Gamma$ be a geodesic. Without loss of generality, we may assume that $t,s\in\mathbb{Z}$. Consider the post composition of $f$ with the left multiplication by $f(s)^{-1}$, which is an isometry on $(\Gamma,|\cdot|)$. Note that the image of $f(s)^{-1}\circ f$ lies in $B_{\Gamma}(e,S)$. Therefore the map
$\mathfrak{o}_{n}\circ f(s)^{-1}\circ f:[0,S^{\prime}]\cap\mathbb{Z}\to X$
is $(\cal C,D+1,L+1)$-Morse. Thus
$\mathfrak{o}_{n}\circ f=\sigma_{n}(f(s))\circ\mathfrak{o}_{n}\circ f(s)^{-1}\circ f$
[…]
Let $B$ denote the ball $B\coloneqq\{g\in G:d(go,o)\leq D^{\prime}\}$. By [3], there exists a uniform constant $Q>0$ such that for all $g\in G$ and $b_{1},b_{2}\in B$,
$\|\mu(g)-\mu(b_{1}gb_{2})\|\leq Q.$ (4.3)
Let $\gamma\in\Gamma$ and $n\geq n_{0}$. We claim that there exists $u\in\cal C$, depending on $\gamma$ and $n$, such that
$\|\mu(\sigma_{n}(\gamma))-u\|\leq Q.$ (4.4)
By Claim A, the point $\sigma_{n}(\gamma)o$ lies within the $D^{\prime}$-neighborhood of some $\cal C^{\prime}$-diamond $\Diamond_{\cal C^{\prime}}(x_{n},y_{n})$, where $x_{n},y_{n}$ are a $\cal C^{\prime}$-regular pair such that $d(x_{n},o)\leq D^{\prime}$ and $d(y_{n},\sigma_{n}(\gamma)o)\leq D^{\prime}$. Let $h_{n}\in G$ be such that $h_{n}o$ is an element of the diamond $\Diamond_{\cal C^{\prime}}(x_{n},y_{n})$ such that $d(\sigma_{n}(\gamma)o,h_{n}o)\leq D^{\prime}$, that is,
$\sigma_{n}(\gamma)=h_{n}b_{n}\quad\text{ for some $b_{n}\in B$.}$
By the description of the diamond, there exists $g_{n}\in G$ such that $x_{n}=g_{n}o$ and $h_{n}o=g_{n}(\exp u)o$ for some $u\in{\cal C}^{\prime}$. Hence
$\sigma_{n}(\gamma)=h_{n}b_{n}=g_{n}(\exp u)k_{n}b_{n}$
for some $k_{n}\in K$.
Since $g_{n}$ and $k_{n}b_{n}$ belong to $B$, we have by (4.3) that
$\|\mu(\sigma_{n}(\gamma))-u\|\leq Q,$
proving the claim (4.4). Therefore for all $n\geq n_{0}$, $\mu(\sigma_{n}(\Gamma))$ is contained in the $Q$-neighborhood of $\cal C^{\prime}$ and hence
$\mathcal{L}_{\sigma_{n}(\Gamma)}\subset\cal C^{\prime}.$
Since $\cal C^{\prime}-\{0\}$ lies in $\cal C_{0}$, we have $\mathcal{L}_{\sigma_{n}(\Gamma)}-\{0\}\subset\cal C_{0}$ for all $n\geq n_{0}$. This finishes the proof of Theorem 4.6. ∎
We now deduce the upper semicontinuity of limit cones from Lemma 4.4 and Proposition 4.5:
Theorem 4.6 (Upper semicontinuity).
Let $\Gamma$ be a $\theta$-Anosov $\theta$-convex subgroup of $G$. If $\sigma_{n}\to\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$, then any accumulation point $\mathcal{L}$ of the sequence $\mathcal{L}_{\sigma_{n}(\Gamma)}$ in $\mathsf{C}(\mathfrak{a}^{+})$ is contained in $\mathcal{L}_{\Gamma}$.
Proof.
As $\Gamma$ is $\theta$-Anosov and $\theta$-convex, $\mathcal{L}_{\Gamma}$ is $\theta$-admissible. Hence the claim follows from Proposition 4.5 by setting $\cal D=\mathcal{L}_{\Gamma}$. ∎
Theorem 1.1 in the introduction is a consequence of Proposition 2.3 and Theorem 4.6.
An application to sharpness
Recall from the introduction that for a closed subgroup $H$ of $G$, a discrete subgroup $\Gamma$ of $G$ is called sharp for $G/H$ if $\mathcal{L}_{\Gamma}\cap\mathcal{L}_{H}=\{0\}$. We show that the sharpness is an open condition in the following situation:
Corollary 4.7.
Let $H<G$ be a closed subgroup. Let $\Gamma$ be a $\theta$-Anosov subgroup of $G$. Suppose that there exists a $\theta$-admissible closed cone $\cal C\subset\mathfrak{a}^{+}$ such that
[…]
Then there exists an open neighborhood $\cal O$ of $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ such that for all $\sigma\in\cal O$, $\sigma(\Gamma)$ is sharp for $G/H$.
In particular, if $H$ is reductive with co-rank one in $G$ and $\Gamma$ is $\theta$-Anosov and sharp for $G/H$, then there exists an open neighborhood $\cal O$ of $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ such that for all $\sigma\in\cal O$, $\sigma(\Gamma)$ is sharp for $G/H$.
Proof.
Since $\cal C$ is a $\theta$-admissible cone such that $\cal C-\{0\}$ is contained in the open cone $\mathfrak{a}^{+}-\mathcal{L}_{H}$ in $\mathfrak{a}^{+}$, by Lemma 4.4, there exists a $\theta$-admissible cone $\cal C_{1}$ in $\mathfrak{a}^{+}$ such that
[…]
Since $\mathcal{L}_{\Gamma}\subset\operatorname{int}\cal C_{1}\cup\{0\}$, by 4.5, we have that for all $\sigma\in\operatorname{Hom}(\Gamma,G)$ sufficiently close to $\operatorname{id}_{\Gamma}$, $\mathcal{L}_{\sigma(\Gamma)}-\{0\}\subset\operatorname{int}\cal C_{1}$, implying that $\mathcal{L}_{\sigma(\Gamma)}\cap\mathcal{L}_{H}=\{0\}$. This proves the first claim.
To prove the second claim, let $H$ be a reductive subgroup of $G$ with co-rank one. Using a compatible Cartan involution, we may assume that we have a Cartan subalgebra $\mathfrak{a}_{H}\subset\mathfrak{a}$ of $H$. Then similarly to the proof of Theorem 2.1, we get
$\mathcal{L}_{H}=\mathcal{W}\mathfrak{a}_{H}\cap\mathfrak{a}^{+}.$
We may assume without loss of generality that $\theta$ is a maximal subset for which $\Gamma$ is $\theta$-Anosov. We claim that
$\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{1}\quad\text{is connected}$
where $\mathfrak{a}^{1}$ is the unit sphere in $\mathfrak{a}$. To see this, note that $\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{1}$ is connected [7, Proposition A.2]. By the maximality assumption on $\theta$, we have that $\mathcal{L}_{\Gamma}\cap\ker\alpha\neq\{0\}$ for every $\alpha\in\Pi-\theta$. In particular, if we denote by $w_{\alpha}$ the reflection along the wall $\ker\alpha$, then for all $\alpha\in\Pi-\theta$,
$(\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{1})\cup w_{\alpha}(\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{1})$
is connected, since $w_{\alpha}$ fixes the intersection $\mathcal{L}_{\Gamma}\cap\ker\alpha\cap\mathfrak{a}^{1}$ pointwise. Since $\mathcal{W}_{\theta}$ is generated by $\{w_{\alpha}:\alpha\in\Pi-\theta\}$, the connectedness of $\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{1}$ follows.
Since $\Gamma$ is sharp for $G/H$, we have $\mathcal{L}_{\Gamma}\cap\mathcal{L}_{H}=\{0\}$. Moreover, since $\mathcal{L}_{\Gamma}-\{0\}$ is connected and $\mathcal{L}_{H}=\mathcal{W}\mathfrak{a}_{H}\cap\mathfrak{a}^{+}$, the set $\mathcal{L}_{\Gamma}-\{0\}$ must lie in the unique connected component, say $\cal C$, of $\mathcal{W}_{\theta}\mathfrak{a}^{+}-\mathcal{W}\mathfrak{a}_{H}$. Note that $\cal C$ is a convex open cone in $\mathcal{W}_{\theta}\mathfrak{a}^{+}$ containing $\mathcal{L}_{\Gamma}-\{0\}$. Since $\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}-\{0\}$ is connected, it follows that
$\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}-\{0\}\subset\cal C.$
For each $w\in\mathcal{W}_{\theta}$, since $w\cal C$ is a connected component of $\mathcal{W}_{\theta}\mathfrak{a}^{+}-\mathcal{W}\mathfrak{a}_{H}$ containing $w(\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}-\{0\})=\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}-\{0\}$, we must have $w\cal C=\cal C$, and hence
$\mathcal{W}_{\theta}\cal C=\cal C.$
In particular, the closure ¯$\bar{\cal}C$ of $\cal C$ in $\mathfrak{a}$ is a $\mathcal{W}_{\theta}$-invariant closed convex cone.
Let
$\cal C_{0}:=\mathfrak{a}^{+}\cap\bar{\cal}C;$
thus, $\mathcal{W}_{\theta}\cal C_{0}=\cal C$. Note that $\operatorname{int}\cal C_{0}$ is a connected component of $\mathfrak{a}^{+}-\mathcal{W}\mathcal{L}_{H}$ containing $\mathcal{L}_{\Gamma}-\{0\}$. Since $\mathcal{L}_{H}$ and $\mathcal{L}_{\Gamma}$ are both $\operatorname{i}$-invariant, it follows that $\cal C_{0}$ is also $\operatorname{i}$-invariant. Let $\cal D$ be any $\theta$-admissible cone containing $\mathcal{L}_{\Gamma}$ (see Lemma 5.4 for the existence of such a cone). Then the cone
$\cal C_{1}:=\cal C_{0}\cap\cal D$
is a $\theta$-admissible cone containing $\mathcal{L}_{\Gamma}-\{0\}$ in its interior, so the second claim follows from the first one. ∎
As mentioned before, the second claim of Corollary 4.7 was obtained in [17, Corollary 6.3] by a different approach.
Examples of $\theta$-Anosov and $\theta$-convex subgroups
Note that since $\mathcal{W}_{\Pi}=\{e\}$, Zariski dense Borel-Anosov subgroups are $\Pi$-convex by Theorem 2.1. Any $\theta$-Anosov subgroup whose limit cone is contained in $\mathfrak{a}_{\theta}^{+}$ with convex limit cone is clearly $\theta$-convex as $\mathcal{W}_{\theta}$ fixes $\mathfrak{a}_{\theta}^{+}$. As another example of $\theta$-Anosov and $\theta$-convex subgroups, we have the following:
Lemma 4.8.
If $L$ is a rank one simple algebraic subgroup of $G$ and $\Gamma$ is a non-elementary convex cocompact subgroup of $L$, then $\Gamma$ is a $\theta$-Anosov $\theta$-convex subgroup of $G$ for some non-empty $\theta\subset\Pi$.
Proof.
We note that Guichard-Wienhard [12, Proposition 4.7] showed that $\Gamma$ is a $\theta$-Anosov subgroup of $G$ for some $\theta$. However, since we also need to demonstrate that $\Gamma$ is $\theta$-convex for an appropriate choice of $\theta$, we provide a complete proof of this result.
If the rank of $G$ is one, then $\Gamma$ is a convex cocompact subgroup of $G$. So the claim just follows since any non-elementary convex cocompact subgroup $\Gamma$ of a rank one Lie group $G$ is $\Pi$-Anosov and $\mathcal{L}_{\Gamma}=\mathfrak{a}^{+}$.
Hence we assume that $\operatorname{rank}G\geq 2$. Since $L<G$ is semisimple, there exists a compatible Cartan involution, and therefore we can assume that we have a Cartan decomposition $L=(K\cap L)(A\cap L)(K\cap L)$ where $K\cap L$ is a maximal compact subgroup and $A\cap L$ is a one-dimensional real split torus. If $A\cap L=\exp(\mathbb{R}v)$ for some $v\in\log(A\cap L)$, then $w\exp(\mathbb{R}_{+}v)w^{-1}\subset A^{+}$ for some Weyl element $w\in\mathcal{W}$. Since $\exp\mathbb{R}_{+}v$ and $\exp\mathbb{R}_{-}v$ are conjugate by an element of $K\cap L$, we have that $\mu(L)=\mathcal{L}_{\Gamma}$ is a ray.
Set
$\theta\coloneqq\{\alpha\in\Pi:\mu(L)\cap\ker\alpha=0\}.$
Since #$\#\Pi\geq 2$ and $\mu(L)$ is a ray, $\theta\neq\emptyset$. Since $\mathcal{L}_{\Gamma}$ is $\operatorname{i}$-invariant, $\theta$ is $\operatorname{i}$-invariant. Moreover, by definition
$\mu(L)\subset\bigcap_{\alpha\in\Pi-\theta}\ker\alpha=\mathfrak{a}_{\theta}.$
Since $\mathcal{W}_{\theta}$ fixes $\mathfrak{a}_{\theta}$ pointwise, $\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}=\mathcal{L}_{\Gamma}$ is a ray and hence convex.
Since $\Gamma<L$ is a convex cocompact subgroup, then $\Gamma$ is quasi-isometrically embedded in $L/(K\cap L)$ under the orbit map $\gamma\mapsto\gamma(K\cap L)$. Since $L/(K\cap L)\subset G/K$ is an isometric embedding, the orbit map $\gamma\mapsto\gamma o$ is a quasi-isometric embedding. Since $\mu(\Gamma)$ lies in a ray $\mu(L)$ which is disjoint from $\bigcup_{\alpha\in\theta}\ker\alpha$ except at zero, it satisfies (3.1) for all $\alpha\in\theta$. Therefore $\Gamma$ is $\theta$-Anosov and $\theta$-convex. ∎
As a third family of examples, we have the following:
Lemma 4.9.
If $\operatorname{rank}G=2$, then any Anosov subgroup of $G$ is $\theta$-convex, where $\theta$ is the maximal subset of $\Pi$ for which $\Gamma$ is $\theta$-Anosov.
Proof.
Since $\mathcal{L}_{\Gamma}\cap\mathfrak{a}^{1}$ is connected [7, Proposition A.2] and $\text{dim}(\mathfrak{a})=2$, $\mathcal{L}_{\Gamma}$ is a convex cone. If $\Gamma$ is Borel-Anosov, the claim follows from the convexity of $\mathcal{L}_{\Gamma}$. Otherwise, it is $\theta$-Anosov with $\theta$ a singleton. By the maximality assumption on $\theta$, $\mathcal{L}_{\Gamma}$ is a convex cone containing the ray $\mathfrak{a}_{\theta}^{+}$. Therefore, $\mathcal{W}_{\theta}\mathcal{L}_{\Gamma}$ is convex. ∎
Therefore we get the following corollary of Theorem 1.1:
Corollary 4.10.
Suppose one of the following:
  1. (1)
    $\Gamma$ is a $\theta$-Anosov subgroup of $G$, such that $\mathcal{L}_{\Gamma}$ is a convex cone contained in $\mathfrak{a}_{\theta}^{+}$ (e.g., a Zariski dense Borel-Anosov subgroup of $G$);
  2. (2)
    $\Gamma$ is any Anosov subgroup of $G$ when $\operatorname{rank}G=2$;
  3. (3)
    $\Gamma$ is a convex cocompact subgroup of a rank-one simple algebraic subgroup $L$ of $G$.
If $\Gamma$ has reducitive Zariski closure, then the map $\sigma\mapsto\mathcal{L}_{\sigma(\Gamma)}$ is continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
As mentioned in the introduction, the third case is due to Kassel [15].
5. Continuity of $\theta$-limit cones
As before, let $G$ be a connected semisimple real algebraic group. We fix a non-empty $\operatorname{i}$-invariant subset $\theta\subset\Pi$. Let
$p_{\theta}:\mathfrak{a}\to\mathfrak{a}_{\theta}$
be the unique projection invariant under $\mathcal{W}_{\theta}$. Set $\mathfrak{a}_{\theta}^{+}=\mathfrak{a}_{\theta}\cap\mathfrak{a}^{+}$. For a discrete subgroup $\Gamma$ of $G$, its $\theta$-limit cone $\mathcal{L}_{\Gamma}^{\theta}$ is defined as
$\mathcal{L}_{\Gamma}^{\theta}:=p_{\theta}(\mathcal{L}_{\Gamma})\subset\mathfrak{a}_{\theta}^{+}.$
We denote by $\mathsf{C}(\mathfrak{a}_{\theta}^{+})$ the space of all closed cones of $\mathfrak{a}_{\theta}^{+}$ equipped with the topology: $\cal C_{i}\to\cal C$ if and only if the Hausdorff distance between $\mathbb{P}(\cal C_{i})$ and $\mathbb{P}(\cal C)$ tends to 00. In this section, we prove the continuity of $\theta$-limit cones $\mathcal{L}_{\Gamma}^{\theta}$ of $\theta$-Anosov subgroups $\Gamma$ of $G$ by a similar argument as the proof of Theorem 1.1.
Theorem 5.1.
Let $\Gamma$ be a $\theta$-Anosov subgroup of $G$ such that $\mathcal{L}_{\Gamma}^{\theta}$ is convex. Then the map
$\sigma\mapsto\mathcal{L}_{\sigma(\Gamma)}^{\theta}\in\mathsf{C}(\mathfrak{a}_{\theta}^{+})$
is continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
Since the projection $p_{\theta}$ is $\mathcal{W}_{\theta}$-invariant, we have that $p_{\theta}^{-1}(\mathcal{L}_{\Gamma}^{\theta})$ is a $\theta$-convex cone for $\Gamma$ as above, and this is why we do not require the $\theta$-convexity in Theorem 5.1.
Remark 5.2.
When $\Gamma$ is a fundamental group of a closed negative curved manifold, embedded in $G$ as a Zariski dense subgroup, Sambarino [30] proved this theorem using thermodynamic formalism. Given the work of Bridgeman, Canary, Labourie and Sambarino [5] which provides the thermodynamic formalism for a general $\theta$-Anosov subgroup, his argument should extend to a general Zariski dense $\theta$-Anosov subgroup.
The purpose of this section is to present an entirely different proof of this result using Proposition 4.5. We believe that this geometric proof is more intuitive and has the potential to extend to a more general setup.
We begin with the following observation using which we will be getting the lower semicontinuity of $\theta$-limit cones without reductive Zariski closure assumption:
Lemma 5.3.
For any $\theta$-Anosov subgroup $\Gamma$ of $G$, we have
$\mathcal{L}_{\Gamma}^{\theta}=p_{\theta}(\mathcal{L}_{\Gamma}^{\rm Jor}).$
Proof.
Let $G_{0}$ be the Zariski closure of $\Gamma$ and let $G_{0}=HU$ be the Levi decomposition where $U$ is the unipotent radical of $G_{0}$ and $H$ is a reductive subgroup. For $g\in G_{0}$, write $g_{H}\in H$ for the $H$-component of $g$. By [34, Lemma 2.11], there exists $C>0$ such that for each $\alpha\in\theta$ and each $\gamma\in\Gamma$,
$\omega_{\alpha}\left(\mu(\gamma)-\mu(\gamma_{H})\right)\leq C$ (5.1)
where $\omega_{\alpha}$ is the fundamental weight corresponding to $\alpha$. Let $\Gamma_{H}$ denote the subgroup $\{\gamma_{H}\in L:\gamma\in\Gamma\}$. This is a discrete subgroup by (5.1) which is Zariski dense in $H$. By Theorem 2.1, we have $\mathcal{L}_{\Gamma_{H}}=\mathcal{L}_{\Gamma_{H}}^{\operatorname{Jor}}$. Since $\lambda(g)=\lambda(g_{H})$ for any $g\in G_{0}$, we have $\mathcal{L}_{\Gamma_{H}}^{\operatorname{Jor}}=\mathcal{L}_{\Gamma}^{\operatorname{Jor}}$. Now, choose a linear form $\psi$ on $\mathfrak{a}_{\theta}$ which is positive on $\mathfrak{a}_{\theta}^{+}-\{0\}$; then it is a non-negative linear combination of $\omega_{\alpha}$, $\alpha\in\theta$. On the other hand, the norm on $\mathfrak{a}_{\theta}^{+}$ is comparable to $\psi$. Therefore it follows from (5.1) that
$\sup_{\gamma\in\Gamma}\|p_{\theta}(\mu(\gamma))-p_{\theta}(\mu(\gamma_{H}))\|<\infty.$
This implies that $\mathcal{L}_{\Gamma}^{\theta}=\mathcal{L}_{\Gamma_{H}}^{\theta}$ and hence $\mathcal{L}_{\Gamma}^{\theta}=p_{\theta}(\mathcal{L}_{\Gamma}^{\operatorname{Jor}})$. ∎
As before, we need to know the existence of $\theta$-admissible cones provided by the following lemma, proved in [8] for $\cal D=\mathcal{L}_{\Gamma}$. The same proof works and we reproduce it for readers’ convenience.
Lemma 5.4.
Let $\cal D\subset\mathfrak{a}^{+}$ be a closed $\operatorname{i}$-invariant cone such that $p_{\theta}(\cal D)$ is convex and $\cal D\cap\ker\alpha=\{0\}$ for all $\alpha\in\theta$. Then for any open cone $\cal C_{0}$ in $\mathfrak{a}^{+}_{\theta}$ containing $p_{\theta}(\cal D)$, there exists a $\theta$-admissible cone $\cal C$ in $\mathfrak{a}^{+}$ such that
$\cal D-\{0\}\subset\operatorname{int}\cal C\quad\text{ and}\quad p_{\theta}(\cal C)\subset\cal C_{0}\cup\{0\}.$
Proof.
Since $p_{\theta}(\cal D)$ is convex and $\operatorname{i}$-invariant, we may assume without loss of generality that $\cal C_{0}$ is convex and $\operatorname{i}$-invariant. Let $H$ be defined as in (4.2) with respect to $\cal D$ and $\cal C_{0}$. Define
[…]
Then $\cal C$ is a closed cone whose interior contains $\cal D-\{0\}$, $\cal C\cap\left(\bigcup_{\alpha\in\theta}\ker\alpha\right)=\{0\}$, $p_{\theta}(\cal C)\subset\cal C_{0}\cup\{0\}$ and $\operatorname{i}(\cal C)=\cal C$. Since $p_{\theta}:\mathfrak{a}\to\mathfrak{a}_{\theta}$ is $\mathcal{W}_{\theta}$-invariant, the preimage $p_{\theta}^{-1}(\cal C_{0})$ is a $\mathcal{W}_{\theta}$-invariant convex cone of $\mathfrak{a}^{+}$. Hence we have
[…]
so $\cal W_{\theta}{}\cal C$ is convex, being the intersection of convex subsets. Therefore $\cal C$ is $\theta$-admissible. ∎
Proof of Theorem 5.1.
The proof of Proposition 2.3 can be easily adapted to show that for any finitely generated subgroup $\Gamma<G$, if $\sigma_{n}$ is a sequence of discrete representations converging to $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$, then any accumulation point $\mathcal{L}$ of $\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}$ in $\mathsf{C}(\mathfrak{a}_{\theta}^{+})$ contains $p_{\theta}(\mathcal{L}_{\Gamma}^{\operatorname{Jor}})$. Therefore the lower semicontinuity follows from Lemma 5.3. To prove the upper semicontinuity, let $\cal C_{0}$ be any open cone in $\mathfrak{a}^{+}_{\theta}$ containing $\mathcal{L}_{\Gamma}^{\theta}-\{0\}$. By applying Lemma 5.4 with $\cal D=\mathcal{L}_{\Gamma}$, we can choose a $\theta$-admissible closed cone $\cal C\subset\mathfrak{a}^{+}$ such that
$\mathcal{L}_{\Gamma}-\{0\}\subset\operatorname{int}\cal C\quad\text{ and}\quad\cal C\subset p_{\theta}^{-1}(\cal C_{0}\cup\{0\}).$
By Proposition 4.5, for all sufficiently large $n$,
$\mathcal{L}_{\sigma_{n}(\Gamma)}\subset p_{\theta}^{-1}(\cal C_{0}\cup\{0\}),$
and hence $\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}\subset\cal C_{0}\cup\{0\}$. This implies the upper semi-continuity, as desired. ∎
We will need the following corollary of the lower semi-continuity of the limit cones in the next section:
Corollary 5.5.
Let $\Gamma$ be a Zariski dense $\theta$-Anosov subgroup, and let $\sigma_{n}\to\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$. For any closed cone $\cal C$ contained in $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}\cup\{0\}$, we have
$\cal C\subset\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}\cup\{0\}\quad\text{ for all large $n$.}$
Proof.
Since the set of Zariski dense representations of $\Gamma$ is Zariski open in $\operatorname{Hom}(\Gamma,G)$ [1, Proposition 8.2] and non-empty since $\Gamma$ is Zariski dense, we may assume without loss of generality that for all $n\geq 1$, $\sigma_{n}(\Gamma)$ is Zariski dense in $G$ , and hence $\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}$ is a convex cone with non-empty interior.
Let $\cal C$ be a closed cone contained in $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}\cup\{0\}$. Let $Z=\cal C\cap\mathfrak{a}^{1}$. Suppose that for some sequence $n_{k}\in\mathbb{N}$ going to $\infty$, we have a sequence of vectors $v_{k}\in Z-\operatorname{int}\mathcal{L}_{\sigma_{n_{k}}(\Gamma)}^{\theta}$. Since $\mathcal{L}^{\theta}_{\sigma_{n_{k}}(\Gamma)}$ is convex, there exists a closed half-space $H_{k}\subset\mathfrak{a}_{\theta}$ whose boundary contains $v_{k}\in Z$ and $\mathcal{L}^{\theta}_{\sigma_{n_{k}}(\Gamma)}\subset H_{k}$.
Since $Z$ is compact, there exists $\varepsilon>0$ such that the $2\varepsilon$-neighborhood of $Z$ is contained in $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$. Let $B_{k}=B(v_{k},2\varepsilon)\subset\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$ denote the ball centered at $v_{k}$ of radius $\varepsilon>0$. Since $v_{k}$ lies in the boundary of $H_{k}$, the $\varepsilon$-neighborhood of $H_{k}$ does not contain $B$ and hence the $\varepsilon$-neighborhood of $\mathcal{L}_{\sigma_{n_{k}}(\Gamma)}\cap\mathfrak{a}^{1}$ does not contain $B_{k}$ and, in particular, it does not contain $\mathcal{L}_{\Gamma}^{\theta}$. Therefore, for any limit $\mathcal{L}$ of the sequence $\mathcal{L}^{\theta}_{\sigma_{n_{k}}(\Gamma)}$ in $\mathsf{C}(\mathfrak{a}_{\theta}^{+})$, the $\varepsilon$-neighborhood of $\mathcal{L}\cap\mathfrak{a}^{1}$ does not contain $\mathcal{L}_{\Gamma}^{\theta}\cap\mathfrak{a}^{1}$. This contradicts the lower semi-continuity part of Theorem 5.1. Hence $Z\subset\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}$ and consequently $\cal C\subset\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}\cup\{0\}$ for all large $n$. ∎
6. $\theta$-growth indicators vary continuously
Let $G$ be a connected semisimple real algebraic group. We fix a non-empty $\operatorname{i}$-invariant subset $\theta\subset\Pi$. In this section, we show that the $\theta$-growth indicators vary continuously in the subspace of Zariski dense $\theta$-Anosov representations of $\operatorname{Hom}(\Gamma,G)$. We consider the $\theta$-Cartan projection:
$\mu_{\theta}=p_{\theta}\circ\mu:G\to\mathfrak{a}_{\theta}^{+}$
Definition 6.1 ($\theta$-growth indicator).
Let $\Gamma$ be a $\theta$-discrete subgroup of $G$, i.e., $\mu_{\theta}|_{\Gamma}$ is a proper map. The $\theta$-growth indicator $\psi_{\Gamma}^{\theta}:\mathfrak{a}_{\theta}^{+}\to\{-\infty\}\cup\mathbb{R}$ is defined as follows: if $u\in\mathfrak{a}^{+}_{\theta}$ is non-zero,
$\psi_{\Gamma}^{\theta}(u)=\|u\|\inf_{u\in\cal C}\tau^{\theta}_{\mathcal{C}}$ (6.1)
where $\cal C\subset\mathfrak{a}^{+}_{\theta}$ ranges over all open cones containing $u$, and $\psi_{\Gamma}^{\theta}(0)=0$. Here $-\infty\leq\tau^{\theta}_{\cal C}\leq\infty$ denotes the abscissa of convergence of the series ${\mathcal{P}}^{\theta}_{\cal C}(s)=\sum_{\gamma\in\Gamma,\mu_{\theta}(\gamma)\in\mathcal{C}}e^{-s\|\mu_{\theta}(\gamma)\|}$.
This definition is equivalent to the one given in (1.1), as was studied in ([28], [19]). The $\theta$-discreteness hypothesis is necessary for the $\theta$-growth indicator to be well-defined. It follows from Definition 3.1 that $\theta$-Anosov subgroups are $\theta$-discrete. For $\theta=\Pi$, we simply write $\psi_{\Gamma}=\psi_{\Gamma}^{\Pi}$. For any discrete subgroup $\Gamma<G$, we have
$\psi_{\Gamma}\leq 2\rho$ (6.2)
where $\rho=\frac{1}{2}\sum_{\alpha\in\Phi^{+}}m(\alpha)\alpha$ is the half sum of all positive roots of $G$ counted with multiplicity (cf. [28]).
The main goal of this section is to prove the following theorem:
Theorem 6.2.
Let $\Gamma$ be a Zariski dense $\theta$-Anosov subgroup of $G$. Let $\sigma_{n}\to\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$. Then for any $v\in\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$, we have
$\lim_{n\to\infty}\psi^{\theta}_{\sigma_{n}(\Gamma)}(v)=\psi^{\theta}_{\Gamma}(v).$ (6.3)
Moreover, the convergence is uniform on compact subsets of $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$.
We begin by recalling the following property of the $\theta$-growth indicator:
Lemma 6.3.
([28], [19]) For any Zariski dense $\theta$-discrete subgroup $\Gamma$ of $G$, $\psi_{\Gamma}^{\theta}$ is upper semicontinuous and concave. Moreover, $\psi_{\Gamma}^{\theta}$ is strictly positive on $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$ and $\mathcal{L}_{\Gamma}^{\theta}=\{\psi_{\Gamma}^{\theta}\geq 0\}$.
As a consequence of these properties, $\psi_{\Gamma}^{\theta}$ is continuous on $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$, since it is a real valued concave function on $\mathcal{L}_{\Gamma}$ and concave functions are continuous on the interior of their domain.
Continuity of $\theta$-limit sets
Let $P_{\theta}$ denote a standard parabolic subgroup of $G$ corresponding to $\theta$; that is, $P_{\theta}$ is generated by the centralizer of $A$ and all root subgroups $U_{\alpha}$, $\alpha\in\Phi^{+}\cup[\Pi-\theta]$ where $[\Pi-\theta]$ denotes the set of all roots in $\Phi$ which are $\mathbb{Z}$-linear combinations of $\Pi-\theta$. We set
$\cal F_{\theta}=G/P_{\theta}.$
Definition 6.4 ($\theta$-limit set).
For a discrete subgroup $\Gamma<G$, the $\theta$-limit set $\Lambda^{\theta}_{\Gamma}$ is a closed subset of $\cal F_{\theta}$ defined as the set of all limit points $\lim k_{i}P_{\theta}\in\cal F_{\theta}$ where $k_{i}\in K$ is a sequence such that there exists $\gamma_{i}=k_{i}a_{i}k_{i}^{\prime}\in\Gamma\cap KA^{+}K$ with $\alpha(\log a_{i})\to\infty$ for all $\alpha\in\theta$.
One important feature of a $\theta$-Anosov subgroup $\Gamma$ is that $\Gamma$ is a hyperbolic group. If $\partial_{\infty}\Gamma$ denotes the Gromov boundary of $\Gamma$, then there exists a $\Gamma$-equivariant continuous injection $f:\partial_{\infty}\Gamma\to\cal F_{\theta}$ such that for $\xi\neq\eta$ in $\partial_{\infty}\Gamma$, $f(\xi)$ and $f(\eta)$ are antipodal, that is, the pair $(f(\xi),f(\eta))$ belongs to the unique open $G$-orbit in $G/P_{\theta}\times G/P_{\theta}$ under the diagonal action, whose image $f(\partial\Gamma)\subset\cal F_{\theta}$ coincides with the $\theta$-limit set $\Lambda^{\theta}_{\Gamma}$.
Let $\mathsf{C}(\cal F_{\theta})$ be the space of all closed subsets of $\cal F_{\theta}$ equipped with the Chabauty topology: $F_{i}\to F$ if and only if any $x\in F$ is a limit of some sequence $x_{i}\in F_{i}$ and the limit of any convergent sequence $x_{i_{j}}\in F_{i_{j}}$ belongs to $F$.
Theorem 6.5.
[12, Theorem 5.13] Let $\Gamma$ be a $\theta$-Anosov subgroup. If $\sigma_{n}\to\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ as $n\to\infty$, then $\Lambda_{\sigma_{n}(\Gamma)}^{\theta}$ converges to $\Lambda_{\Gamma}^{\theta}$ in $\mathsf{C}(\cal F_{\theta})$.
Conformal measures on limit sets
The space $\mathfrak{a}_{\theta}^{*}$ of all linear forms on $\mathfrak{a}_{\theta}$ can be identified with the following space:
$\mathfrak{a}_{\theta}^{*}=\{\phi\in\mathfrak{a}^{*}:\phi=\phi\circ p_{\theta}\}.$
In particular, we may regard a linear form on $\mathfrak{a}_{\theta}$ as a linear form on the whole space $\mathfrak{a}$.
Definition 6.6.
For $\phi\in\mathfrak{a}_{\theta}^{*}$, a probability measure $\nu$ on $\cal F_{\theta}$ is called a $(\Gamma,\phi)$-conformal measure if for all $\gamma\in\Gamma$ and $\xi\in\cal F_{\theta}$,
$\frac{d\gamma_{*}\nu}{d\nu}(\xi)=e^{\phi(\beta_{\xi}(o,\gamma o))}$ (6.4)
where $\beta$ denotes the $\mathfrak{a}$-valued Busemann map ([27]).
A linear form $\phi\in\mathfrak{a}_{\theta}^{*}$ is said to be tangent to $\psi_{\Gamma}^{\theta}$ at a non-zero vector $v\in\mathfrak{a}_{\theta}^{+}$ if
$\phi\geq\psi_{\Gamma}^{\theta}\quad$ and $\quad\phi(v)=\psi_{\Gamma}^{\theta}(v)$.
The following gives a complete classification of conformal measures supported on the $\theta$-limit sets:
Theorem 6.7.
([22], [31], [19, Corollary 1.13]) Let $\Gamma$ be a Zariski dense $\theta$-Anosov subgroup. For any unit vector $v\in\operatorname{int}\mathcal{L}^{\theta}_{\Gamma}$, there exists a unique tangent linear form $\phi_{v}\in\mathfrak{a}_{\theta}^{*}$ to $\psi_{\Gamma}^{\theta}$ at $v$ and there exists a unique $(\Gamma,\phi_{v})$-conformal measure $\nu_{v}$ on $\Lambda_{\Gamma}^{\theta}$. Moreover, the assignment $v\mapsto\phi_{v}\mapsto\nu_{v}$ gives bijections among the set of all unit vectors $\operatorname{int}\mathcal{L}^{\theta}_{\Gamma}$, the set of all tangent linear forms to $\psi_{\Gamma}^{\theta}$ and the set of all $\Gamma$-conformal measures supported on $\Lambda_{\Gamma}^{\theta}$.
Continuous variations of growth indicators
If $\phi\in\mathfrak{a}_{\theta}^{*}$ is positive on $\mathcal{L}_{\Gamma}^{\theta}-\{0\}$, the critical exponent $0<\delta_{\phi,\Gamma}<\infty$, which is the abscissa of convergence of the series $s\mapsto\sum_{\gamma\in\Gamma}e^{-s\phi(\mu(\gamma))}$, is well-defined [19, Lemma 4.2] and is equal to
$\delta_{\phi,\Gamma}^{\theta}\coloneqq\limsup_{T\to\infty}\frac{1}{T}{\log\#\{\gamma\in\Gamma:\phi(\mu(\gamma))\leq T\}}.$ (6.5)
Theorem 6.8.
Let $\Gamma$ be a $\theta$-Anosov subgroup with $\mathcal{L}_{\Gamma}^{\theta}$ convex. Let $\phi\in\mathfrak{a}_{\theta}^{*}$ be positive on $\mathcal{L}_{\Gamma}^{\theta}-\{0\}$. Then for all $\sigma\in\operatorname{Hom}(\Gamma,G)$ sufficiently close to $\operatorname{id}_{\Gamma}$, the critical exponent $0<\delta_{\phi,\sigma(\Gamma)}<\infty$ is well-defined and the map
$\sigma\mapsto\delta_{\phi,\sigma(\Gamma)}$
is continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
Proof.
In [5, Proposition 8.1] (see also [6, Section 4.4]), this is obtained for any $\phi\in\mathfrak{a}_{\theta}^{*}$ which is positive on $\mathfrak{a}_{\theta}^{+}-\{0\}$, but the condition $\phi>0$ on $\mathfrak{a}_{\theta}^{+}-\{0\}$ was needed only to guarantee that $\phi$ is positive on $\mathcal{L}_{\sigma(\Gamma)}^{\theta}$ for all $\sigma$ sufficiently close to $\operatorname{id}_{\Gamma}$. By Theorem 5.1, if $\phi>0$ on $\mathcal{L}_{\Gamma}^{\theta}-\{0\}$, then $\phi>0$ on $\mathcal{L}_{\sigma(\Gamma)}^{\theta}-\{0\}$ for all $\sigma$ sufficiently close to $\operatorname{id}_{\Gamma}$ and therefore the above formulation holds. ∎
We now give a proof of Theorem 6.2; the use of conformal measures in this context was inspired by the work of Sullivan [32] and McMullen [24] where conformal measures were used to study critical exponents of Kleinian groups.
Proof of Theorem 6.2 Let $v\in\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$ be a unit vector. By Theorem 6.7, there exists a unique linear form $\phi=\phi_{v}\in\mathfrak{a}_{\theta}^{*}$ tangent to $\psi_{\Gamma}^{\theta}$ at $v$. The $\theta$-Anosov property of $\Gamma$ implies that $\psi_{\Gamma}^{\theta}$ is vertically tangent, i.e., there is no linear form tangent to $\psi_{\Gamma}^{\theta}$ at a non-zero vector $v\in\partial\mathcal{L}_{\Gamma}^{\theta}$ ([31], [19, Theorem 12.2]). This implies that $\phi>0$ on $\mathcal{L}_{\Gamma}^{\theta}-\{0\}$. By Theorem 6.8, we have $0<\delta_{\phi,\sigma_{n}(\Gamma)}<\infty$ for all large $n$. This implies that $\delta_{\phi,\sigma_{n}(\Gamma)}\phi$ is a tangent linear form [19, Lemma 4.5], in particular,
$\psi^{\theta}_{\sigma_{n}(\Gamma)}\leq\delta_{\phi,\sigma_{n}(\Gamma)}\phi.$ (6.6)
Since $\delta_{\phi,\sigma_{n}(\Gamma)}\to\delta_{\phi,\Gamma}$ by Theorem 6.8 and $\delta_{\phi,\Gamma}=1$ [19, Corollary 4.6], we have
lim $\limsup\psi^{\theta}_{\sigma_{n}(\Gamma)}(v)\leq\limsup\delta_{\phi,\sigma_{n}(\Gamma)}\phi(v)=\phi(v)=\psi^{\theta}_{\Gamma}(v).$ (6.7)
By 5.5,
$v\in\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}$
for all large $n$. Therefore, by Theorem 6.7, there exists a linear form $\phi_{n}\in\mathfrak{a}_{\theta}^{*}$ tangent to $\psi_{\sigma_{n}(\Gamma)}^{\theta}$ at $v$ and a $(\sigma_{n}(\Gamma),\phi_{n})$-conformal probability measure, say $\nu_{n}$, on the limit set $\Lambda_{\sigma_{n}(\Gamma)}^{\theta}\subset\cal F_{\theta}$.
We claim that by passing to a subsequence, $\nu_{n}$ weakly converges to a $\Gamma$-conformal measure on $\Lambda_{\Gamma}^{\theta}$. To see this, fix a closed cone $\cal D$ such that $v\in\operatorname{int}\cal D\subset\cal D\subset\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}\cup\{0\}$. By Corollary 5.5, $\cal D\subset\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}\cup\{0\}$ for all large enough $n\geq 1$. Since $\psi_{\sigma_{n}(\Gamma)}^{\theta}\geq 0$ on $\mathcal{L}^{\theta}_{\sigma_{n}(\Gamma)}$, we have that for all large $n$,
$\phi_{n}\geq 0\quad\text{ on $\cal D$. }$
Moreover, since $\cal D$ has non-empty interior and $\phi_{n}(v)=\psi_{\sigma_{n}(\Gamma)}^{\theta}(v)\leq\psi_{\Gamma}^{\theta}(v)$ for all large $n$ by (6.7), it follows that the sequence $\phi_{n}$ converges, up to passing to a subsequence, to some linear form, say $\phi\in\mathfrak{a}_{\theta}^{*}$, which can possibly be the zero linear form at this point. Let $\nu$ be a weak-limit of the sequence $\nu_{n}$. Since $\sigma_{n}\to\operatorname{id}_{\Gamma}$, it follows from (6.4) that $\nu$ is a $(\Gamma,\phi)$-conformal measure on $\cal F_{\theta}$. Since $\Lambda_{\sigma_{n}(\Gamma)}^{\theta}$ converges to $\Lambda_{\Gamma}^{\theta}$ as $n\to\infty$ by Theorem 6.5, $\nu$ is indeed supported on $\Lambda_{\Gamma}^{\theta}$. This proves the claim.
Therefore by Theorem 6.7, we have
$\phi\geq\psi_{\Gamma}^{\theta};$
for instance, $\phi$ cannot be zero. Hence
$\psi_{\Gamma}^{\theta}(v)\leq\phi(v)=\lim\phi_{n}(v)=\lim\psi^{\theta}_{\sigma_{n}(\Gamma)}(v).$ (6.8)
Together with (6.7), this proves
$\psi_{\Gamma}^{\theta}(v)=\lim\psi^{\theta}_{\sigma_{n}(\Gamma)}(v).$
Finally, for any compact subset $Z\subset\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$, we need to show that as $n\to\infty$,
$\psi^{\theta}_{\sigma_{n}(\Gamma)}|_{Z}\to\psi^{\theta}_{\Gamma}|_{Z}\quad\text{uniformly on }Z.$ (6.9)
By 5.5, we have $Z\subset\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}$ for all sufficiently large $n\geq 1$. Since $\operatorname{int}\mathcal{L}^{\theta}_{\Gamma}$ is a convex subset, any compact subset $Z\subset\operatorname{int}\mathcal{L}^{\theta}_{\Gamma}$ can be covered by finitely many compact convex subsets of $\operatorname{int}\mathcal{L}^{\theta}_{\Gamma}$. Hence we may assume without loss of generality that $Z$ is a compact convex subset.
Recall that each $\psi^{\theta}_{\sigma_{n}(\Gamma)}$ is continuous on $\operatorname{int}_{\sigma_{n}(\Gamma)}^{\theta}$ (see the remark following Lemma 6.3), and hence we may assume that $\psi^{\theta}_{\sigma_{n}(\Gamma)}|_{Z}$ is continuous for all $n\geq 1$. We claim that the family $\cal E_{Z}\coloneqq\{\psi^{\theta}_{\sigma_{n}(\Gamma)}|_{Z}:n\in\mathbb{N}\}$ is equicontinuous. Observe that $\cal E_{Z}$ consists of positive concave continuous functions. Moreover, by [19, Lemma 3.13],
$\psi_{\sigma_{n}(\Gamma)}^{\theta}(u)\leq\max_{v\in p_{\theta}^{-1}(u)}\psi_{\sigma_{n}(\Gamma)}(v)\quad\text{for all }u\in\mathfrak{a}^{+}_{\theta}.$
Since $\psi_{\sigma_{n}(\Gamma)}\leq 2\rho$ by (6.2), we have that
$\max_{u\in Z}\psi_{\sigma_{n}(\Gamma)}^{\theta}(u)\leq\max_{v\in p_{\theta}^{-1}(Z)\cap\mathcal{L}_{\sigma_{n}(\Gamma)}}2\rho(v).$ (6.10)
Consider $\theta$-admissible closed cones $\cal D\subset\cal C$ such that $\mathcal{L}_{\Gamma}-\{0\}\subset\operatorname{int}\cal D$ and $\cal D-\{0\}\subset\operatorname{int}\cal C$, which exist by Lemma 5.4. By 4.5, we have that $\mathcal{L}_{\sigma_{n}(\Gamma)}\subset\cal C$ for all large $n\geq 1$. Since $\cal C$ is $\theta$-admissible and hence $\cal C\cap\ker\alpha=\{0\}$ for all $\alpha\in\theta$, the intersection $p^{-1}_{\theta}(Z)\cap\cal C$ is compact. Thus the right side of (6.10) is uniformly bounded by
$\max_{v\in p_{\theta}^{-1}(Z)\cap\cal C}2\rho(v)<\infty.$
Hence, $\cal E_{Z}$ is a uniformly bounded family of concave functions. Since this is true for an arbitrary compact convex subset $Z\subset\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$, it follows that the functions in $\cal E_{Z}$ are uniformly Lipschitz, which implies that they are equicontinuous.
By the Arzelà–Ascoli theorem, any sequence from the family $\cal E_{Z}$ has a subsequence that converges uniformly. Since the growth indicator functions $\psi^{\theta}_{\sigma_{n}(\Gamma)}|_{Z}$ converge pointwise to $\psi_{\Gamma}^{\theta}|_{Z}$, any uniform limit of $\cal E_{Z}$ must be $\psi_{\Gamma}^{\theta}|_{Z}$. This proves (6.9), completing the proof of Theorem 6.2.
Proof of Corollary 1.5: Let $u_{\Gamma}^{\theta}\in\mathfrak{a}_{\theta}$ be the unique unit vector defined by
$\delta^{\theta}_{\Gamma}=\psi^{\theta}_{\Gamma}(u_{\Gamma}^{\theta})=\max_{\|u\|=1,u\in\mathfrak{a}_{\theta}^{+}}\psi^{\theta}_{\Gamma}(u).$ (6.11)
As $\Gamma$ is $\theta$-Anosov, we have that $\psi_{\Gamma}^{\theta}$ is strictly concave and vertically tangent [19, Theorem 12.2]. It follows that $u_{\Gamma}^{\theta}\in\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}.$
Choose a compact neighborhood $B\subset\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}\cap\mathfrak{a}^{1}_{\theta}$ of $u_{\Gamma}^{\theta}$. Let $\sigma_{n}$ be a sequence converging to $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$. By 5.5, we have
$B\subset\operatorname{int}\mathcal{L}_{\sigma_{n}(\Gamma)}^{\theta}\cap\mathfrak{a}^{1}_{\theta}\quad\text{for all sufficiently large }n\geq 1.$
Let $u_{n}\in B$ be the unit vector such that $\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{n})$ attains the maximum of $\psi_{\sigma_{n}(\Gamma)}^{\theta}$ on $B$. We claim that
$u_{n}=u^{\theta}_{\sigma_{n}(\Gamma)},$ (6.12)
that is, $\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{n})=\max_{v\in\mathfrak{a}^{1}_{\theta}}\psi_{\sigma_{n}(\Gamma)}^{\theta}(v)$.
By definition of $u_{\Gamma}^{\theta}$, there exist $\eta>0$ such that
$\psi_{\Gamma}^{\theta}(u_{\Gamma}^{\theta})\geq\eta+\max_{v\in\partial B}\psi_{\Gamma}^{\theta}(v).$
Since $\psi_{\sigma_{n}(\Gamma)}^{\theta}$ converges uniformly to $\psi_{\Gamma}^{\theta}$ on $B$ by Theorem 6.2, we have, for all sufficiently large enough $n\geq 1$,
$\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{\Gamma}^{\theta})\geq\frac{\eta}{2}+\max_{v\in\partial B}\psi_{\sigma_{n}(\Gamma)}^{\theta}(v).$ (6.13)
Since $\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{n})\geq\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{\Gamma}^{\theta})$, it follows from (6.13) that $u_{n}$ lies in in the interior of $B$ for all large $n\geq 1$. Now suppose that $u_{n}\neq u^{\theta}_{\sigma_{n}(\Gamma)}$. This means that $u^{\theta}_{\sigma_{n}(\Gamma)}\notin B$. Noting that both $u_{n}$ and $u^{\theta}_{\sigma_{n}(\Gamma)}$ lie in $\operatorname{int}\mathcal{L}^{\theta}_{\Gamma}$, which is convex, consider the line segment in $\operatorname{int}\mathcal{L}_{\Gamma}^{\theta}$ connecting $u_{n}$ to $u^{\theta}_{\sigma_{n}(\Gamma)}$:
$t\mapsto c(t)\coloneqq tu_{n}+(1-t)u^{\theta}_{\sigma_{n}(\Gamma)}\quad\text{for }t\in[0,1].$
Note that $10<\|c(t)\|<$ 1 for all $t\in(0,1)$. Therefore, by the concavity and homogeneity of $\psi_{\sigma_{n}(\Gamma)}^{\theta}$, we have that for all $t\in(0,1)$,
$\psi_{\sigma_{n}(\Gamma)}^{\theta}\left(\frac{c(t)}{\|c(t)\|}\right)$ $=\frac{1}{\|c(t)\|}\psi_{\sigma_{n}(\Gamma)}^{\theta}(c(t))$
$>\psi_{\sigma_{n}(\Gamma)}^{\theta}(c(t))$
$\geq t\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{n})+(1-t)\psi_{\sigma_{n}(\Gamma)}^{\theta}(u^{\theta}_{\sigma_{n}(\Gamma)})$
$\geq\min\left\{\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{n}),\,\psi_{\sigma_{n}(\Gamma)}^{\theta}(u^{\theta}_{\sigma_{n}(\Gamma)})\right\}$
$=\psi_{\sigma_{n}(\Gamma)}^{\theta}(u_{n}).$
Since $\frac{c(t)}{\|c(t)\|}\to u_{n}$ as $t\to 1$, this is a contradiction to the fact that $u_{n}$ attains the maximum of $\psi_{\sigma_{n}(\Gamma)}^{\theta}$ on $B$. Therefore, $u^{\theta}_{\sigma_{n}(\Gamma)}\in B$, proving the claim (6.12).
Since $B$ is an arbitrary compact neighborhood of $u_{\Gamma}^{\theta}$, it follows from (6.12) that
$u_{n}=u^{\theta}_{\sigma_{n}(\Gamma)}\to u^{\theta}_{\Gamma}\quad\text{as }n\to\infty.$ (6.14)
Since $\psi_{\sigma_{n}(\Gamma)}^{\theta}\to\psi_{\Gamma}^{\theta}$ uniformly on $B$, by Theorem 6.2, we obtain from (6.14) that
$\delta_{\sigma_{n}(\Gamma)}^{\theta}\to\delta_{\Gamma}^{\theta}\quad\text{ as $n\to\infty$.}$
This finishes the proof.
7. Continuity of more general critical exponents
If $\Gamma$ is a discrete subgroup of $G$ and $\phi\in\mathfrak{a}^{*}$ is positive on $\mathcal{L}_{\Gamma}-\{0\}$, then the $\phi$-critical exponent $\delta_{\phi,\Gamma}$ of $\Gamma$ is well-defined as in (6.5).
In this section, we obtain the following continuity of $\phi$-critical exponents as an application of Theorem 4.5. The main point of this theorem is that we are not assuming that the linear form $\phi$ is $p_{\theta}$-invariant unlike in Theorem 6.8.
Theorem 7.1 (Theorem 1.6).
Let $\Gamma$ be a $\theta$-Anosov subgroup of $G$ with $\mathcal{L}_{\Gamma}\subset\mathfrak{a}_{\theta}^{+}$ convex. Let $\phi\in\mathfrak{a}^{*}$ be such that $\phi>0$ on $\mathcal{L}_{\Gamma}-\{0\}$. Then for all $\sigma\in\operatorname{Hom}(\Gamma,G)$ sufficiently close to $\operatorname{id}_{\Gamma}$, the critical exponent $0<\delta_{\phi,\sigma(\Gamma)}<\infty$ is well-defined and the map $\sigma\mapsto\delta_{\phi,\sigma(\Gamma)}$ is continuous at $\operatorname{id}_{\Gamma}\in\operatorname{Hom}(\Gamma,G)$.
Proof.
Let $\phi\in\mathfrak{a}^{*}$ be such that $\phi>0$ on $\mathcal{L}_{\Gamma}-\{0\}$. Set
$\phi^{\prime}\coloneqq\phi\circ p_{\theta},$
so that $\phi^{\prime}\in\mathfrak{a}_{\theta}^{*}$. Since $\phi^{\prime}=\phi$ on $\mathfrak{a}_{\theta}$ and $\mathcal{L}_{\Gamma}=\mathcal{L}_{\Gamma}^{\theta}$, we have $\phi^{\prime}>0$ on $\mathcal{L}_{\Gamma}-\{0\}$. Fix a small open neighborhood $\cal O^{\prime}$ of $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ so that
$c\coloneqq\sup_{\sigma\in\cal O^{\prime}}\delta_{\phi^{\prime},\sigma(\Gamma)}<\infty$
by Theorem 6.8. For $\eta>0$, set ${\mathcal{N}}_{\eta}(\mathcal{L}_{\Gamma})\coloneqq\{v\in\mathfrak{a}^{+}:\|v-\mathcal{L}_{\Gamma}\|\leq\eta\|v\|\}.$
[…]
$-\varepsilon\phi(v)\leq(\phi-\phi^{\prime})(v)\leq\varepsilon\phi(v).$ (7.1)
We can take a small neighborhood $\cal O\subset\cal O^{\prime}$ of $\operatorname{id}_{\Gamma}$ in $\operatorname{Hom}(\Gamma,G)$ so that for any $\sigma\in\cal O$, $\sigma$ is $\theta$-Anosov and
$\mathcal{L}_{\sigma(\Gamma)}\subset{\mathcal{N}}_{\eta}(\mathcal{L}_{\Gamma})$ (7.2)
by Theorem 4.6. In particular, there exists a finite subset $F_{\sigma}\subset\Gamma$ such that
$\mu(\sigma(\Gamma-F_{\sigma}))\subset{\mathcal{N}}_{\eta}(\mathcal{L}_{\Gamma}).$
Since $\lim\delta_{\phi^{\prime},\sigma(\Gamma_{0})}=\delta_{\phi^{\prime},\Gamma_{0}}$ as $\sigma\to\operatorname{id}_{\Gamma}$ by Theorem 6.8, we may assume that for all $\sigma\in\cal O$, we have
$|\delta_{\phi^{\prime},\sigma(\Gamma)}-\delta_{\phi^{\prime},\Gamma}|\leq\varepsilon$ (7.3)
by replacing $\cal O$ by a smaller one if necessary. Then for any $\sigma\in\cal O$, we have that for all $s>0$,
$\sum_{\gamma\in\Gamma-F_{\sigma}}e^{-(1-\varepsilon)s\phi(\mu(\sigma(\gamma)))}\geq\sum_{\gamma\in\Gamma-F_{\sigma}}e^{-s\phi^{\prime}(\mu(\sigma(\gamma)))}.$
It follows that
$\delta_{(1-\varepsilon)\phi,\sigma(\Gamma)}\geq\delta_{\phi^{\prime},\sigma(\Gamma)}\text{ and hence }\delta_{\phi,\sigma(\Gamma)}\geq(1-\varepsilon)\delta_{\phi^{\prime},\sigma(\Gamma)}.$
Similarly, we have that for any $\sigma\in\cal O$,
$\sum_{\gamma\in\Gamma-F_{\sigma}}e^{-(1+\varepsilon)s\phi(\mu(\sigma(\gamma)))}\leq\sum_{\gamma\in\Gamma-F_{\sigma}}e^{-s\phi^{\prime}(\mu(\sigma(\gamma)))}.$
Therefore
$\delta_{(1+\varepsilon)\phi,\sigma(\Gamma)}\leq\delta_{\phi^{\prime},\sigma(\Gamma)}\text{ and hence }\delta_{\phi,\sigma(\Gamma)}\leq(1+\varepsilon)\delta_{\phi^{\prime},\sigma(\Gamma)}.$
Therefore for all $\sigma\in\cal O$,
$(1-\varepsilon)\delta_{\phi^{\prime},\sigma(\Gamma)}\leq\delta_{\phi,\sigma(\Gamma)}\leq(1+\varepsilon)\delta_{\phi^{\prime},\sigma(\Gamma)}.$
It follows from (7.3) that for all $\sigma\in\cal O$,
$|\delta_{\phi,\sigma(\Gamma)}-\delta_{\phi,\Gamma}|\leq 2\varepsilon\delta_{\phi^{\prime},\sigma(\Gamma)}\leq 2c\varepsilon.$
This finishes the proof. ∎
Proof of Corollary 1.7: Since $\phi>0$ on $\mathcal{L}_{\Gamma}-\{0\}$, we have that $\phi>0$ on $\mathcal{L}_{\sigma_{n}(\Gamma)}-\{0\}$ for all sufficiently large $n$ by Theorem 4.6. It follows from [18] that $\delta_{\phi,\sigma_{n}(\Gamma)}$ is the minimum constant $\kappa_{n}\geq 0$ such that
$\psi_{\sigma_{n}(\Gamma)}\leq\kappa_{n}\phi.$
Therefore $\kappa\geq\delta_{\phi,\sigma_{n}(\Gamma)}$. By Theorem 1.6, we have
$\kappa\geq\delta_{\phi,\Gamma}.$
This implies that $\psi_{\Gamma}\leq\kappa\phi$, as desired.